### v3

# A Step to Decouple Optimization in 3DGS

Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for real-time novel view synthesis. As an explicit representation optimized through gradient propagation among primitives, optimization widely accepted in deep neural networks (DNNs) is actually adopted in 3DGS, such as synchronous weight updating and Adam with the adaptive gradient. However, considering the physical significance and specific design in 3DGS, there are two overlooked details in the optimization of 3DGS: (i) update step coupling, which induces optimizer state rescaling and costly attribute updates outside the viewpoints, and (ii) gradient coupling in the moment, which may lead to under- or over-effective regularization. Nevertheless, such a complex coupling is under-explored. After revisiting the optimization of 3DGS, we take a step to decouple it and recompose the process into: Sparse Adam, Re-State Regularization and Decoupled Attribute Regularization. Taking a large number of experiments under the 3DGS and 3DGS-MCMC frameworks, our work provides a deeper understanding of these components. Finally, based on the empirical analysis, we re-design the optimization and propose AdamW-GS by re-coupling the beneficial components, under which better optimization efficiency and representation effectiveness are achieved simultaneously.
1 Introduction
Novel view synthesis is a fundamental task in both computer graphics and computer vision. Owing to its highly parallelized design and efficient use of GPU, 3DGS (Kerbl et al., 2023) leverages explicit primitive-based representations to achieve significantly higher efficiency than implicit representation NeRF (Mildenhall et al., 2020), while still delivering competitive reconstruction quality. 3DGS sparks a wave of research and rapid extensions to various downstream applications (Matsuki et al., 2024; Chen et al., 2024; Dai et al., 2024).
As a representation directly optimized via backpropagation, 3DGS inherits the Adam (Kinga et al., 2015) from DNNs, and in default attributes of all primitives are updated simultaneously within each iteration, which we call synchronous optimization in this work. Given the viewpoint relationship, an ideal scenario is that the optimization of primitives is guided by gradients from visible viewpoints in different directions. However, upon closer examination, we observe that even when primitives are invisible under a given viewpoint and their gradients are set to zero, the update-step coupling induced by Adam and synchronous optimization implicitly rescales the optimizer state and causes updating in these invisible viewpoints. From the perspective of optimization efficiency, (Mallick et al., 2024) introduces Sparse Adam, which restricts updates to primitives visible under viewpoints while filtering out those invisible with zero gradients, thereby enabling asynchronous optimization. Nevertheless, this optimizer has not been widely adopted, since its efficiency gains come at the expense of degraded performance. However, we think that such a viewpoint-related optimizer remains promising, with the current limitation lying in the insufficient recognition of update-step coupling.
Furthermore, previous study (Loshchilov & Hutter, 2017) in DNNs points out the gradient coupling in Adam makes the regularization unstable. In current 3DGS, a variety of regularization losses (Yu et al., 2024) are introduced to improve reconstruction quality or mitigate redundancy111Redundancy is strictly defined as minimizing active primitives without sacrificing reconstruction quality.. With limited regularization intensity controlled via adaptive gradients and decoupled from the photometric loss, a natural question arises as to whether existing regularization is appropriately effective. In practice, the regularization loss is thought to be controlled through hyperparameters, yet it remains unclear whether they provide sufficient flexibility. Moreover, can further decoupling enable more precise control over regularization itself? In this work, we focus on opacity and scaling regularization, which are closely tied to reconstruction quality in the 3DGS-MCMC framework (Kheradmand et al., 2024) and redundancy removal in the vanilla 3DGS framework (Papantonakis et al., 2024). Prior studies have shown that the densification stage often produces a large number of redundant primitives (Liu et al., 2025b), while the framework lacks a native mechanism to automatically remove them. Even with opacity regularization, existing pipelines usually rely on additional pruning operations or direct modifications to the densification process. However, if regularization can be well recomposed, is it possible to achieve redundancy removal without resorting to such extra pruning operations?
Considering the coupling in 3DGS optimization, this work takes a step to decouple it and recompose it. Building on the analysis, the optimization is then redesigned to improve efficiency and/or reconstruction quality. To summarize, our contributions are:
Reanalyzing the complex coupling during 3DGS optimization, we take a step to decouple it and recompose it into 3 effective components: Sparse Adam, Re-State Regularization and Decoupled Attribute Regularization.
We take a step to explore the contribution of them, including the characteristic of Sparse Adam, activation of regularization, role of the regularization hyperparameter, risk of implicit update and necessity of controllable regularization as well as exploration.
Based on the experiments and analysis, we propose the AdamW-GS with controllable attribute regularization and adopt it into vanilla 3DGS, 3DGS-MCMC and more variants. Experiments show that our optimization method yields better reconstruction quality and optimization efficiency, and significantly reduces redundancy in 3DGS without introducing additional pruning operations.
2 Related Work
“Coupling” in 3DGS Optimization: 3DGS commonly inherits the Adam optimizer from DNNs. However, its synchronous optimization causes primitives in invisible viewpoints to undergo overlooked update steps. Sparse Adam (Mallick et al., 2024) introduces asynchronous updates that improve efficiency, yet the insufficient recognition of update-step coupling results in degraded performance. Regularization in 3DGS is also coupled with the photometric loss, akin to L2 weight regularization in adaptive gradient algorithms for DNNs. As analyzed in AdamW (Loshchilov & Hutter, 2017), the coupling of adaptive gradients in Adam makes the effective strength of regularization governed by the second moment, thereby motivating weight decay decoupling. A similar idea is reflected in (Rota Bulò et al., 2025), which—though not explicitly designed for decoupling—replaces opacity reset in vanilla 3DGS with a constant opacity decay during densification, effectively constructing an AdamW-style optimizer. Nevertheless, this approach neglects the physical properties of opacity and the varying significance of individual primitives. The central contribution of this work is to further decouple and recompose 3DGS optimization and, through this process, reinterpret redundancy in 3DGS and reconstruction quality in 3DGS-MCMC from an optimization perspective.
Redundancy during Optimization: Typically, 3DGS requires careful treatment of the special process of primitive generation and pruning (or death (Kheradmand et al., 2024; Zhu et al., 2025)). Current studies generally observe that optimization produces abundant redundant primitives, while pruning (in adaptive density control (Kerbl et al., 2023)) alone remains insufficient to eliminate them. Four main strategies have been proposed to address this redundancy: (1) Densification refinement (Fang & Wang, 2024; Mallick et al., 2024; Wang et al., 2025); (2) Hand-crafted criterion (Fan et al., 2024; Niemeyer et al., 2025; Hanson et al., 2025); (3) Learning-based pruning (Lee et al., 2024; Liu et al., 2025b; Zhang et al., 2024a). More details can be found in Sec. H. (4) Optimization-related operations: Optimization together with additional opacity L1 regularization has become a common technique for redundancy removal, even extending beyond Efficient 3DGS tasks (Papantonakis et al., 2024; Lee et al., 2024; Kheradmand et al., 2024; Liu et al., 2025a; Svitov et al., 2024).
3 Preliminary: 3DGS(-MCMC) and Adam
3DGS approximates the radiance field of a target scene using a set of Gaussian primitives parameterized by location, opacity (), scale (), and other attributes, and renders images from given viewpoints via alpha blending. Further details are provided in Appendix Sec. D.
Typically, the training pipeline can be divided into densification and pure optimization (P-Op). During densification, new primitives are generated to enhance scene representation while redundant ones are pruned (low-opacity primitives are removed) (Kerbl et al., 2023), or new primitives are sampled from existing primitives while dead primitives are reallocated to new locations (Kheradmand et al., 2024). After densification, there is a P-Op stage, during which only gradient propagation and attribute updates occur. In this training pipeline, photometric loss (left in Eq.1) is adopted with Adam optimizer. For 3DGS-MCMC, extra regularization loss in Eq. 1 is utilized to promote respawning.
In this work, and denote photometric loss and any regularization, while and are the corresponding gradient.
Adam is the first-order stochastic optimization method that adaptively estimates moment statistics of the gradients. Specifically, it maintains exponential moving averages of the first moment and second moment , which are bias-corrected to and . The parameter update is then performed as Eq.3, where denotes the learning rate and ensures numerical stability.
We omit the expression of bias correction procedure; in this paper, and denote the primitive attribute in and gradient information respectively.
Consistent with the DNNs(Goodfellow et al., 2016), Adam by default updates all attributes synchronously, including both the primitives visible under the current viewpoint and those that are not. In spite of the 0 gradient, they are still included in the update step, during which their moment are rescaled and attributes are still updated. In this paper, updates of points in invisible viewpoints are referred to as Implicit Update, which are influenced by the previous moment.
4 Methodology
4.1 Characteristics of Sparse Adam
| Cite | PSNR | SSIM | LPIPS | /m | /m |
| GS1 | 27.507 | 0.815 | 0.216 | 3.331 | 0.232 |
| GS2 | 27.285 | 0.809 | 0.228 | 2.532 | 0.039 |
| GS3 | 27.567 | 0.816 | 0.216 | 3.342 | 0.048 |
As discussed in Sec.1, using Sparse Adam directly improves the optimization efficiency but causes degraded performance. Denoted as a filter that takes the value only for primitives visible under the current viewpoint and for others, Sparse Adam can be obtained by replacing in Eq. 2 via:
However, from the perspective of inter-viewpoint updates, Sparse Adam is effectively viewpoint-stable: the update of primitives is influenced only by steps in visible viewpoints at the moment and does not undergo implicit updates in invisible viewpoints.
To conduct a preliminary investigation, three experiments were performed on vanilla 3DGS: (GS1) training with the original Adam optimizer, (GS2) replacing Adam with Sparse Adam, and (GS3) employing Adam during the densification stage while applying Sparse Adam in the P-Op stage. Based on these comparisons in Table 1, we make the following observations: Observation 1 Sparse Adam is more stable: Since densification is governed by gradient magnitudes, the smaller number of primitives observed under Sparse Adam suggests that more primitives fail to meet the gradient threshold. Moreover, Sparse Adam holds on more active primitives, especially the GS3 in Table 1 (0.048 million dead primitives in Sparse Adam v.s. 0.232 million in Adam), which implies components in Adam are potential for pruning redundant primitives. Observation 2 Sparse Adam is less explorative: although it quickly drives primitives toward stability and results in fewer primitives overall, its performance degrades noticeably. Experiments on 3DGS-MCMC show the similar observation in some scenes that Sparse Adam makes smaller reallocated primitive number, and this phenomenon can be found in Figure 3 f.
4.2 Update-step decoupling: Re-State Regularization
The key question is what components differentiate Adam from Sparse Adam. To answer this, we revisit synchronous optimization with Adam introduced in Sec.3. The ignored update-step coupling arises as zero-gradient updates rescale the moment and subsequently induce implicit update based on the rescaled moment. This first implies that the optimizer state of primitives in invisible viewpoints continues to change. As visualized in Figure 1 a, the second moment in 3DGS-MCMC with Adam is notably smaller than with Sparse Adam, suggesting potential amplification to the first moment. To verify this, we compare the effective strength of opacity regularization in Figure 1 b, , and observe that Adam induces stronger regularization, consistent with our discussion in Sec. 4.1. The moment rescaling helps regularization activation. The comparison in Figure 1 a also shows that the magnitude remains consistently high for Sparse Adam, underscoring the importance of moment rescaling: when inappropriate gradients accumulate in one view, they are difficult to dissipate, thereby hindering optimization. The role of implicit updates is direct: in invisible viewpoints, primitives are continuously updated based on the past moment. However, no prior work has explicitly examined the impact of these components on optimization. Motivated by this, we construct 2 decoupled variants for Sparse Adam.
The implementation of moment rescaling is straightforward. We define an optimization interval and, proposing a milestone style State Sampling Schedule (StSS), uniformly sample the current primitives at the fixed interval. The sampled primitives are directly processed as in Eq. 5 to deliberately attenuate the state. This process is referred to as Re-State Regularization (RSR).
Here, the subscript follows an asynchronous mode different from synchronous mode in Eq.2, meaning that each primitive has its own distinct . A detailed hyperparameters discussion is in Sec. J.2.
We also construct an implicit update to study its role. While implicit updates can be beneficial in some cases, they may introduce negative effects, potentially linked to redundancy (see Appendix Sec. I.4). In the experiment section of the mainbody, we retain implicit updates only as a tool to amplify attribute regularization under Sparse Adam to study the influence of regularization: when in Eq. 6 is large, maintaining implicit updates further strengthens regularization. Concretely, our Artificial Implicit Update (AIU) is defined as a uniform sampler that selects a random subset of invisible primitives for extra updates, while their moments remain unchanged in invisible viewpoints as in Sparse Adam.
where the and the are the scale of image pixel and the primitive number at the present viewpoint respectively, decided by the code project222https://github.com/DerThomy/3dgs-mcmc.
4.3 Gradient Decoupling: Decoupled Attribute Regularization
We view attribute regularization as an essential component of optimization. In efficient 3DGS and related tasks, the widely-used opacity L1 regularization exemplifies this idea, analogous to weight regularization in DNNs, though with a different role. As discussed in Sec. 3, our understanding of how regularization effectively constrains each primitive remains limited. Implicit updates may further exaggerate the effect of regularization: when in Eq. 6 approaches or exceeds , the regularization term continues to influence the optimization via the adaptive gradient or implicit updates. The coupling of photometric and regularization losses in Eq. 6 therefore raises a natural question: is the current regularization over- or under-effective?
First, we return to the factors actually govern the effect of the current attribute regularization loss (L1 loss with the gradient coupling). The role of the hyperparameters serves to maintain stability between losses, ensuring that the regularization remains relatively small at critical moments without interfering with the photometric loss. Relevant results are presented in Appendix Sec. I.1. Regarding the modulation by the moment, the regularization is activated when takes the majority or the moment is rescaled during synchronous optimization in Adam or reset in 3DGS-MCMC.
Using Sparse Adam or Adam in 3DGS-MCMC clearly alters the effect of attribute regularization. As illustrated in Figure 3 e–f, different optimizers yield different numbers of reallocated primitives and PSNR, indicating over- or under-effective regularization. This suggests that stronger regularization may be required to improve reconstruction quality in some cases. However, due to the gradient coupling in the moment (Eq. 6), directly amplifying easily leads to update steps that become overly attribute-dependent and less guided by the photometric gradient, while the coupling also sustains the effect of regularization across iterations. In extreme cases, such as when hyperparameters are scaled by 10×, optimization fails entirely. When amplification relies on the second moment, the effect of regularization depends on its relative scale to . Furthermore, if aligns with , the second moment grows rapidly; if they are opposed, it grows slowly, which is contrary to expectations. This coupling prevents effective control of regularization. As shown in Figure 1 c,d, the order-of-magnitude gap between average and maximum further illustrates that improper amplification under coupling easily results in over-effective regularization. Decoupling and and recomposing the regularization enables more reliable control without destabilizing optimization. Further empirical experiments in Sec. 5 will show the side effect of coupling as well.
Therefore, the first step is to decouple the gradients in the moment (Eq. 6). This step is the same as AdamW, and readily yields Eq. 7, which defines the moment update term containing only the gradient of the photometric loss . For , AdamW constructs updates that are not controlled by the second moment. This AdamW-style decoupling is effectively equivalent to a constant penalty on opacity (Rota Bulò et al., 2025). The experiments are presented in Sec. I.2. Clearly, this approach is not optimal. Unlike DNNs, each attribute in 3DGS has physical meaning, and each primitive carries a distinct importance. Penalizing all primitives with the same strength fails to remove redundant primitives when the penalty is too weak, while overly strong penalties hinder optimization.
AdamW-style decoupling suggests that effective regularization assigns different penalties to different primitives. Such assignment, however, cannot be arbitrary: (i) it must not interfere with normal optimization, particularly in under-constructed regions ( is large); (ii) it must still fulfill its core function of suppressing overfitting when is small (Srivastava et al., 2014). This implies that regularization should adapt to the geometry of the data (Pascanu & Bengio, 2013), naturally linking back to adaptive gradient (Duchi et al., 2011; Kinga et al., 2015) where adjustments are made via . The empirical success of vanilla regularization over the constant penalty further highlights its potential. Since we have already disentangled the moment in Eq. 7, now provides a more faithful estimate of the parameter space. Motivated by this, we introduce the regularization form ( is from Eq. 7). This design brings 4 benefits: (1) updates are preserved in under-constructed regions with large ; (2) when a primitive lies near a saddle point (small ), regularization facilitates escape, consistent with saddle-point analyses in (Wang et al., 2025) and long-axis primitive gradients in (Zhang et al., 2024b); (3) regularization remains small in general but amplifies adaptively when it becomes influential; and (4) it enables explicit control over regularization via moment modulation, e.g., through RSR. Building on this rationale, we extend this approach and propose Eq. 8.
We maintain the loss balance through , and normalize the regularization scale by . Regularization aligned with improves stability while removing dependence on primitive number, which can vary by orders of magnitude across training, views, and scenes. keeps on the same order as maximum step size of Adam, matches the maximum observed in Figure 1 c–d, and extensive results confirm its robustness. Then, this recomposed regularization is referred to as Decoupled Attribute Regularization (DAR). We provide detailed hyperparameter selection in Appendix Sec. J.2.
4.4 Recoupling: AdamW-GS
Building on Sec. 4.1, 4.2, and 4.3, we propose AdamW-GS for better 3DGS optimization by recoupling Sparse Adam, RSR, and DAR. In detail, attributes of primitives are optimized asynchronously with Sparse Adam and supervised via photometric loss and DAR. At fixed training intervals, RSR uniformly samples primitives and rescales their moments via Eq. 5 to better activate regularization. The rescaled moment replaces the old one and directly participates in subsequent optimization.
We consider two DAR variants—opacity and scaling regularization—both of which are already employed in 3DGS-MCMC. Considering the bad transportation in vanilla 3DGS (Jung et al., 2024), exploration in this work means better primitive movement. We directly use noise regularization in 3DGS-MCMC as an extra position regularization, which is similar to DAR, but controlled by opacity and primitive shape not the second moment. However considering the sensitivity to noise (Jung et al., 2024), the noise is excluded for MipNerf360 indoors scenes with fewer primitives. To explore the effectiveness of exploration, we provide contrast experiments with or without opacity reset(Kerbl et al., 2023). Opacity reset encourages exploration via lowering the opacity. Opacity Correction is adopted for cloning in densification as well, which is thought to be necessary (Kheradmand et al., 2024). We adopt the similar operations for opacity as (Rota Bulò et al., 2025).
5 Experiments
Dataset and Metric Following existing research (Kerbl et al., 2023), we employ 13 scenes from 3 datasets, including 9 scenes in Mip-Nerf360 (Barron et al., 2022), 2 scenes from the Tanks&Temples dataset (Knapitsch et al., 2017) and 2 scenes provided by Hedman et al. (Hedman et al., 2018).
We use a train/test split for datasets, using the methodology suggested by Mip-NeRF360, taking every 8th photo for test, for consistent and meaningful comparisons to generate the error metrics, using the standard PSNR, SSIM, and L-PIPS(vgg) metrics used most frequently in the literature. We also present the total primitive number , the active primitives number and the dead primitives number . Given the significant variation in primitive number across methods (often differing by millions), we employ normalized changes to the vanilla (, where ) to better illustrate primitive number variations. The active/dead primitive is defined by opacity larger than 1/255 or not. This definition follows the underlying 3DGS333https://github.com/graphdeco-inria/gaussian-splatting implementation, where primitives with opacity below 1/255 are excluded from rendering, differing from (Kheradmand et al., 2024).
Baselines To evaluate the impact of different components on the training pipeline, we conduct extensive experiments on both the vanilla 3DGS and 3DGS-MCMC frameworks. These two baselines respectively represent approaches without or with an maximum. Without explicitly introducing pruning operations, we observe that a substantial number of redundant primitives can be automatically removed under the vanilla 3DGS with AdamW-GS. For comparison, we include two adaptive pruning methods, learning-based MaskGaussian (Liu et al., 2025b) and Redundant Primitive Removal (RePR) (Papantonakis et al., 2024) with hand-crafted criterion. Unlike approaches with a predefined pruning rate, these methods automatically identify redundant primitives and claim to achieve this with negligible performance degradation. All methods are run on a single A6000 GPU.
| Cite | Sparse Adam | AIU | RSR | PSNR | SSIM | LPIPS | /m | ||||
| 3DGS-MCMC | MC1 | x | x | x | L1 | L1 | 27.948 | 0.833 | 0.199 | -3.75% | 3.313 |
| MC2 | ✓ | x | x | L1 | L1 | 27.998 | 0.832 | 0.199 | +4.28% | 3.313 | |
| MC3 | ✓ | ✓ | x | L1 | L1 | 28.050 | 0.833 | 0.198 | +3.62% | 3.313 | |
| MC4 | ✓ | ✓ | ✓ | L1 | L1 | 28.017 | 0.834 | 0.191 | +0.51% | 3.313 | |
| MC19 | ✓ | x | ✓ | L1 | L1 | 28.075 | 0.837 | 0.190 | +2.97% | 3.313 | |
| MC7 | ✓ | x | ✓ | DAR | DAR | 28.185 | 0.839 | 0.182 | +4.72% | 3.313 | |
| MC8 | ✓ | x | ✓ | DAR | DAR | 28.219 | 0.840 | 0.182 | +4.52% | 3.313 |
5.1 Results and Analysis
Over- or Under-Effective Regularization To examine the controllability of regularization, we introduce AIU during densification as a direct tool to amplify its effect. Its role is straightforward: when the moment of a primitive places it under regularization, AIU preserves its update, yielding more dead primitives and thus making stronger regularization. Figure 3 e–f visualizes the reallocated primitives change in 3DGS-MCMC with AIU. Enhanced regularization promotes exploration and improves reconstruction quality in some cases (MC3 and MC4 in Table 2), but the effect is scene-dependent. For instance, in Kitchen (shown in Figure 3 f), rapid growth of reallocated primitives degrades performance. This suggests that regularization can be under-effective in some scenes, requiring amplification (e.g., via AIU or other methods), but also over-effective when its strength is excessive or inapposite.
Activated via the Moment Within the 3DGS-MCMC framework, we evaluate the effect of our proposed RSR on regularization, a way of activating regularization via rescaling the moment. Regardless of the original L1 form subject to gradient coupling or our proposed DAR, attribute regularization remains tied to the second moment. Figure 3 e–f shows that applying RSR substantially increases the number of dead primitives, thereby strengthening attribute regularization. This directly demonstrates the effectiveness of RSR as a component to amplify regularization. However, as highlighted by the comparison between MCMC-Sparse-RSR and MCMC-AdamWGS-MC17 in Figure 3 e, the influence of RSR with same StSS is weaker under L1 regularization with gradient coupling than under DAR. By decoupling the gradient in the moment and recomposing the attribute regularization, RSR more effectively amplifies its impact, which also translates into improved reconstruction quality under stronger regularization.
Side Effect in Coupling To further examine this effect, we increase the sampling ratio of RSR in MidNeRF-360 indoor scenes (which typically involve fewer primitives), with results summarized in Appendix, from Sec. K.6 to Sec. K.9, MC20 and MC21, in which we replace the low-ratio StSSMC1 in original 3DGSMCMC with L1 regularization and 3DGSMCMC with our proposed AdamW-GS with higher-ratio StSSMC2. As the sampling ratio increases, original 3DGS-MCMC with L1 regularization exhibits significant drops in reconstruction quality across all four scenes (e.g., Room: 32.514→31.179 dB; Kitchen: 32.289→31.924 dB). The same phenomenon is also observed in the experiments on outdoor scenes. In contrast, 3DGS-MCMC with DAR shows little to no degradation in Room, Counter, and Bosai, and only a minor drop in Kitchen (32.546→32.298 dB), while consistently outperforming L1 regularization under the same settings. These results confirm that coupling hinders effective control of regularization and can easily introduce negative effects on reconstruction quality. Thus, decoupling the gradient from the moment and recomposing regularization is necessary for stable and effective optimization.
Experiments with RSR also highlight the issue of over- or under-effective regularization. For scenes with more primitives, such as MipNeRF360 outdoor datasets, we adopt a higher sampling ratio in RSR. As the sampling ratio increases, the number of reallocated primitives grows further (Figure 3 e), strengthening regularization and enhancing the exploration ability of 3DGS-MCMC. A direct comparison between MC7 and MC8 shows that higher sampling ratios improve reconstruction quality. This further demonstrates the effectiveness of RSR to control regularization.
AdamW-GS in 3DGS-MCMC With properly tuned hyperparameters, we evaluate our proposed AdamW-GS within the 3DGS-MCMC framework. As shown in Table 3, our approach achieves superior PSNR, SSIM, and LPIPS over vanilla 3DGS-MCMC. By decoupling gradient coupling and recomposing regularization, we obtain more stable regularization; and by decoupling update-step coupling, our proposed RSR provides finer control of regularization. Furthermore, eliminating numerous ineffective zero-gradient updates via Sparse Adam improves overall optimization efficiency.
| Methods | All | Outdoor | Indoor | ||||||||||
| PSNR | SSIM | LPIPS | time/mins | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||||
| Original MCMC | 27.948 | 0.833 | 0.199 | -3.75% | 46.81 | 25.105 | 0.755 | 0.212 | -4.38% | 31.502 | 0.930 | 0.182 | -2.97% |
| +AdamW-GS(MC8) | 28.219 | 0.840 | 0.182 | +4.52% | 39.77 | 25.247 | 0.764 | 0.191 | +3.07% | 31.934 | 0.935 | 0.172 | +6.33% |
| vanilla 3DGS | 27.506 | 0.815 | 0.216 | (3.098m) | 30.58 | 24.648 | 0.728 | 0.239 | (4.512m) | 31.080 | 0.925 | 0.189 | (1.331m) |
| +AdamW-GS(GS8) | 27.678 | 0.822 | 0.220 | -49.3% | 18.53 | 24.854 | 0.740 | 0.243 | -48.4% | 31.209 | 0.925 | 0.191 | -50.4% |
| +AdamW-GS(GS7) | 27.730 | 0.820 | 0.222 | -46.9% | 19.18 | 24.949 | 0.737 | 0.248 | -44.0% | ||||
| RePR | 27.503 | 0.815 | 0.218 | -41.1% | 26.73 | 24.661 | 0.728 | 0.241 | -41.5% | 31.055 | 0.924 | 0.190 | -40.6% |
| MaskGaussian | 27.485 | 0.815 | 0.219 | -53.1% | 26.80 | 24.683 | 0.728 | 0.240 | -46.4% | 30.988 | 0.924 | 0.192 | -61.5% |
Robust Hyperparameter Test and Autonomously Redundancy Removal Given that AdamW-GS, equipped with DAR and RSR, introduces many hyperparameters, we test their robustness by directly applying the method to vanilla 3DGS with same hyperparameters. Following the training division of vanilla 3DGS, we design several StSS configurations, as shown in Figure 9. Similar to our observations in 3DGS-MCMC with AdamW-GS, where moment rescaling and opacity DAR encourage a large number of dead primitives for exploration, our experiments show that with our proposed AdamW-GS, a large number of redundant primitives can be autonomously removed for 3DGS without extra pruning components, and the improvement in reconstruction quality also indicates enhanced exploration (Observation 2 issue in Sparse Adam in Sec.4.1).
We compare vanilla 3DGS trained with AdamW-GS against two adaptive pruning methods. Importantly, our approach contains no extra pruning component; instead, it relies solely on opacity DAR. In terms of pruning performance, AdamW-GS achieves results comparable to MaskGaussian and even outperforms it by 2% on outdoor scenes. Beyond pruning efficiency, our method also improves reconstruction quality: in outdoor scenes, it reduces the number of primitives by 48.4% while increasing PSNR by 0.2 dB and SSIM by 0.01. In indoor scenes, it removes 50% of primitives while still improving PSNR by 0.1 dB. By contrast, MaskGaussian removes 61% of primitives but degrades PSNR by 0.1 dB, which is because MaskGaussian suffers from potential reconstruction quality risk (see Figure 2 c). Visualization in Figure 3 c-d also demonstrates better detail reconstruction in our methods compared to vanilla 3DGS, which is one of the reasons for relatively more primitives than MaskGaussian. Ultimately, our objective is not simply to minimize the number of active primitives, but to strike a preferable balance—achieving higher reconstruction quality with as few primitives as necessary.
| Deep blending | Tank & Temples | |||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||
| 3DGSMCMC | 30.089 | 0.914 | 0.239 | -24% | 24.563 | 0.869 | 0.160 | -3.1% |
| +AdamW-GS | 30.417 | 0.916 | 0.228 | 2.8% | 24.726 | 0.875 | 0.150 | 6.7% |
| 3DGS | 29.694 | 0.904 | 0.247 | 2.60m | 23.677 | 0.848 | 0.178 | 1.60m |
| +AdamW-GS | 30.260 | 0.912 | 0.245 | -60% | 24.303 | 0.855 | 0.181 | -40% |
| MaskGaussian | 29.895 | 0.908 | 0.248 | -65% | 23.607 | 0.846 | 0.181 | -50% |
Additional datasets are also evaluated with our proposed method. As shown in Table 4, experiments on Deep Blending and Tanks&Temples further demonstrate the superiority of our approach. On Deep Blending in particular, 3DGS equipped with AdamW-GS even surpasses the PSNR of the original 3DGS-MCMC. In Appendix Sec. E.4, we also report results on long-sequence datasets for vanilla 3DGS and 3DGS-MCMC, both with and without our proposed AdamW-GS, where our method continues to show clear advantages.
| Indoor | Outdoor | |||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||
| MaskGaussian | 30.988 | 0.924 | 0.192 | -61.5% | 24.683 | 0.728 | 0.240 | -46.4% |
| +AdamW-GS | 31.199 | 0.925 | 0.193 | -68.6% | 24.939 | 0.739 | 0.244 | -48.1% |
To further validate our findings and the robustness of our method, we apply AdamW-GS to additional pipelines. As shown in Table 3 and discussed therein, MaskGaussian exhibits a potential reconstruction-quality risk on indoor datasets—a phenomenon consistent with Observation 1 in Sec. 4.1 from the Adam vs. Sparse Adam experiments. We attribute this issue to the synchronous updating of mask scores in MaskGaussian (a brief introduction to MaskGaussian is given in Sec. E). Results in Table 5 further support this hypothesis: when trained with AdamW-GS, MaskGaussian no longer suffers from the quality risk and additionally prunes about 7% redundant primitives on indoor scenes. This also demonstrates that our method is compatible with additional pruning operations. In Appendix Sec. E, we provide results of more pipeline variants with AdamW-GS, offering further evidence of the robustness.
To better understand how our method penalizes redundant primitives, we further visualize the primitive number dynamics during training, as shown in Figure 2. PePR, which relies solely on opacity L1 regularization during densification, shows limit in reducing redundancy. In contrast, with our proposed regularization, the number of primitives remains consistently lower. Even in cases where is larger than that of MaskGaussian, our method (vanilla 3DGS with AdamW-GS)—restricted to the densification—still outperforms MaskGaussian in reducing redundancy during densification.
Extra Exploration is Necessary Previous studies have highlighted the role of exploration in 3DGS (Jung et al., 2024; Kheradmand et al., 2024), and here we investigate its interplay with redundancy. AdamW-GS benefits from the noise-based position regularization, yielding improved pruning efficiency and higher SSIM. As shown in Table 2 GS7 and GS8, removing opacity reset weakens the noise-based regularization, leading to less improvement in SSIM and redundancy removal. We attribute the effectiveness of the noise-based regularization in AdamW-GS to its more appropriate attribute regularization. Overall, our results underscore that additional exploration is necessary. However, the noise-based position regularization from 3DGS-MCMC cannot be directly applied to scenes with relatively few primitives (e.g., indoor scenes), where their higher sensitivity to noise results in a drop in reconstruction quality. In Appendix Sec. F, we provide a more detailed analysis of the exploration ability induced by our method. We also introduce two additional exploration strategies that can be combined with current AdamW-GS and yield further improvements on certain indoor scenes as well.
Efficiency Analysis Benefiting from Sparse Adam, as shown in Table 3, 3DGS-MCMC further improves optimization efficiency. Vanilla 3DGS enjoys the same benefit, while our method achieves additional gains by substantially reducing redundancy during the densification stage, thereby further enhancing optimization efficiency. More detailed analysis is provided in Appendix Sec. C.
6 Discussion and Conclusion
This work takes a step to decouple the complex 3DGS optimization and analyzes key mechanisms. AdamW-GS is further proposed, which improves efficiency and enables more controllable regularization. The method enhances reconstruction quality in the MCMC framework and achieves notable redundancy removal in vanilla 3DGS with slight quality gains and even without pruning. Our results indicate that efficiency and effectiveness can be jointly improved given appropriate regularization. While we focused on opacity and scaling, more common DAR in different attribute or alternative strategies may yield further improvements. Despite the robust current hyperparameters and hand-crafted SStS, changing the parameters during training in an adaptive manner may offer a more promising direction. Finally, although noise-driven exploration shows necessary, it can harm certain scenes. Other provided exploration strategies offer limited improvement. A more systematic investigation of exploration strategies may be a promising direction for future research.
Acknowledgments This work was supported by China Mobile Hunan Company Limited, China Mobile Communications Group Co., Ltd., and by the National Natural Science Foundation of China under Grant U22B2050, 62425305, 62221002, and 62503161. This work was conducted as part of the project “Networked Robotic System for Major Equipment Manufacturing (5G+Robotics)”.
Appendix A Reproducibility
The datasets are directly available from (Kerbl et al., 2023). Methodological steps and formulas are detailed in Secs. 4.4 and J.1, while all hyperparameters are reported in Secs. J.2 and K.
More rendering visualization can be found in Figure 10.
Appendix B The Use of Large Language Models (LLMs)
In this work, LLMs are used solely to assist in writing and polishing the manuscript; they do not contribute to the discovery or experimental design.
Appendix C Time Cost Analysis
Within the 3DGS-MCMC framework, employing Sparse Adam leads to a substantial reduction in step time—approximately 50% (see Table 6)—by directly eliminating implicit updates. The fluctuations in forward and backward costs stem from variations in the number of activated primitives during training. Although DAR encourages a larger proportion of dead primitives, Sparse Adam exhibits greater stability and tends to retain more activated primitives.
For vanilla 3DGS, the reduction in time cost benefits from both the decrease in update-step overhead achieved by Sparse Adam and the reduced forward and backward costs resulting from fewer primitives. Overall, our method achieves more than a 40% reduction in total runtime. In indoor scenes with fewer primitives, the update-step cost becomes negligible, as illustrated in Figure 4. Due to extra loss in MaskGaussian as its pruning strategy, it yields almost no improvement in the backward time cost.
| Methods | Avg. | Bicycle | Flowers | Garden | Stump | Treehill | Room | Counter | Kichen | Bosai | |
| vanilla 3DGS | Densification | 0.37 | 0.43 | 0.38 | 0.49 | 0.39 | 0.37 | 0.32 | 0.34 | 0.34 | 0.32 |
| Forward | 6.91 | 8.35 | 6.09 | 8.47 | 6.44 | 6.08 | 6.70 | 6.67 | 7.61 | 5.86 | |
| Backward | 13.72 | 14.26 | 11.27 | 15.75 | 11.54 | 11.38 | 14.69 | 14.62 | 17.38 | 12.59 | |
| Step | 9.55 | 16.30 | 10.25 | 17.08 | 13.65 | 10.53 | 4.63 | 3.75 | 5.69 | 4.12 | |
| All | 30.58 | 39.36 | 28.01 | 41.81 | 32.04 | 28.38 | 26.35 | 25.40 | 31.03 | 22.91 | |
| RePR | Densification | 0.95 | 1.40 | 0.99 | 1.29 | 1.54 | 1.07 | 0.58 | 0.60 | 0.54 | 0.62 |
| Forward | 6.07 | 6.55 | 5.41 | 6.85 | 5.73 | 5.40 | 6.00 | 6.23 | 7.04 | 5.50 | |
| Backward | 12.75 | 12.23 | 10.52 | 13.82 | 10.73 | 10.69 | 13.72 | 14.24 | 16.79 | 12.05 | |
| Step | 6.92 | 11.20 | 7.39 | 11.51 | 11.05 | 7.84 | 3.10 | 2.75 | 4.51 | 3.01 | |
| All | 26.73 | 31.40 | 24.33 | 33.48 | 29.06 | 25.02 | 23.41 | 23.83 | 28.90 | 21.20 | |
| MaskGaussian | Densification | 0.37 | 0.46 | 0.36 | 0.49 | 0.41 | 0.38 | 0.31 | 0.30 | 0.32 | 0.33 |
| Forward | 5.91 | 7.42 | 5.26 | 7.40 | 5.46 | 5.39 | 5.30 | 5.60 | 6.41 | 4.96 | |
| Backward | 13.66 | 14.19 | 11.2 | 15.66 | 11.20 | 11.45 | 14.15 | 14.81 | 17.61 | 12.69 | |
| Step | 6.83 | 12.50 | 7.64 | 12.52 | 9.38 | 8.12 | 2.39 | 2.29 | 3.94 | 2.74 | |
| All | 26.8 | 34.59 | 24.48 | 36.09 | 26.46 | 25.36 | 22.16 | 23.01 | 28.30 | 20.75 | |
| Densification | 0.33 | 0.34 | 0.33 | 0.34 | 0.35 | 0.33 | 0.32 | 0.31 | 0.34 | 0.33 | |
| 3DGS | Forward | 5.19 | 5.21 | 4.45 | 4.81 | 4.75 | 4.49 | 5.54 | 5.69 | 6.27 | 5.56 |
| Backward | 10.80 | 9.76 | 8.59 | 9.70 | 9.07 | 8.61 | 12.1 | 12.73 | 14.25 | 12.47 | |
| +AdamWGS(GS8) | Step | 2.17 | 3.90 | 2.89 | 3.76 | 4.19 | 3.05 | 0.34 | 0.40 | 0.58 | 0.43 |
| All | 18.53 | 19.23 | 16.28 | 18.62 | 18.37 | 16.50 | 18.39 | 19.15 | 21.45 | 18.81 | |
| 3DGSMCMC | Densification | 0.02 | 0.04 | 0.03 | 0.04 | 0.04 | 0.03 | 0.02 | 0.02 | 0.02 | 0.02 |
| Forward | 9.03 | 9.13 | 8.15 | 9.79 | 8.02 | 7.72 | 9.67 | 9.47 | 10.45 | 8.87 | |
| Backward | 18.95 | 17.39 | 15.49 | 19.19 | 14.85 | 14.39 | 22.55 | 22.20 | 24.54 | 19.97 | |
| Step | 21.09 | 31.51 | 19.45 | 33.76 | 24.68 | 20.46 | 10.23 | 8.12 | 12.17 | 8.75 | |
| All | 46.81 | 58.08 | 43.13 | 62.79 | 47.60 | 42.61 | 42.49 | 39.81 | 47.19 | 37.63 | |
| Densification | 0.03 | 0.05 | 0.03 | 0.05 | 0.04 | 0.03 | 0.02 | 0.02 | 0.02 | 0.02 | |
| 3DGSMCMC | Forward | 9.18 | 9.50 | 8.10 | 10.15 | 8.51 | 7.91 | 9.55 | 9.26 | 10.84 | 8.85 |
| Backward | 11.46 | 18.33 | 15.25 | 21.32 | 16.02 | 14.41 | 21.18 | 19.91 | 25.61 | 19.66 | |
| +AdamWGS(MC8) | Step | 11.46 | 18.75 | 12.04 | 20.64 | 14.84 | 12.35 | 6.20 | 5.06 | 7.89 | 5.40 |
| All | 39.77 | 46.65 | 35.44 | 52.17 | 39.41 | 34.71 | 36.97 | 34.27 | 44.39 | 33.94 |
Appendix D Preliminary
3DGS(-MCMC)
3DGS approximates the radiance field of targeted scenes via a group of Gaussian primitives parameterized with where , denotes the primitive position, is the positive semi-definite covariance matrix decoupled into quaternion and a scaling vector, is considered as the opacity value. The color attribute is typically stored as spherical harmonics coefficients and converted into RGB when rendering.
Given a viewpoint and transformation (Kerbl et al., 2023), the pixel on the screen is rendered via -blending according to the projected depth of primitives:
Here the superscript denotes the corresponding parameters under the given viewpoint after the respective transformation, whereas .
Optimization
The 3DGS training pipeline typically initializes from Structure-from-Motion (SfM) (Schonberger & Frahm, 2016) followed by several warm-up iterations (Kerbl et al., 2023; Jung et al., 2024). The primitives then undergo densification, wherein new primitives are generated to enhance scene representation while redundant ones are pruned. In the vanilla pipeline (Figure 9 a), both operations are governed by adaptive density control: new primitives are created via cloning or splitting based on gradient magnitude, whereas low-opacity primitives are removed. After densification, optimization proceeds in a pure optimization (P-Op) stage, during which only gradient propagation and attribute updates occur. Throughout warm-up, densification, and P-Op, training employs a photometric loss (left in Eq. 1) with the Adam optimizer. Building on Stochastic Gradient Langevin Dynamics (SGLD), 3DGS-MCMC (Kheradmand et al., 2024) follows a similar stage division, as shown in Figure 9 b: new primitive sampling and dead primitive reallocation during densification, followed by a shortened P-Op. It also adopts Adam, augmented with a regularization loss in Eq. 1 to promote respawn and additional positional noise for exploration or adherence to the SGLD form.
Appendix E More Pipeline Variants with AdamW-GS or Experiments on Different Datasets
| Pipeline | AdamW-GS | All | Outdoor | Indoor | |||||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||
| vanilla 3DGS | x | 27.506 | 0.815 | 0.216 | (3.098m) | 24.648 | 0.728 | 0.239 | (4.512m) | 31.080 | 0.925 | 0.189 | (1.331m) |
| ✓ | 27.678 | 0.822 | 0.220 | -49.3% | 24.854 | 0.740 | 0.243 | -48.4% | 31.209 | 0.925 | 0.191 | -50.4% | |
| MaskGaussian | x | 27.485 | 0.815 | 0.219 | -53.1% | 24.683 | 0.728 | 0.240 | -46.4% | 30.988 | 0.924 | 0.192 | -61.5% |
| ✓ | 27.721 | 0.821 | 0.221 | -57.2% | 24.939 | 0.739 | 0.244 | -48.1% | 31.199 | 0.925 | 0.193 | -68.6% | |
E.1 MaskGaussian with AdamW-GS: More Stable Updating of Mask Score
MaskGaussian introduces a probabilistic formulation of 3D Gaussian primitives, where each primitive is assigned a learnable probability of existence that governs a dynamic sampling process during rendering. This mechanism enables adaptive pruning of redundant primitives throughout the 3DGS optimization. Specifically, each primitive is associated with a mask score , from which a binary mask is stochastically sampled using the Gumbel-Softmax reparameterization. The standard rendering equation in Eq. 9 is thus modified to incorporate the mask in both color accumulation and transmittance updates, as shown in Eq. 10. Detailed derivations and implementation are provided in (Liu et al., 2025b).
When MaskGaussian is equipped with AdamW-GS, the observed PSNR improvement aligns with that of vanilla 3DGS using AdamW-GS. Additionally, for indoor scenes, it enables an extra pruning of approximately 7% of primitives. This suggests that synchronous updates of mask scores may lead to potentially “destructive pruning behavior”. In contrast, our method cannot only be jointly used with such pruning strategies to stabilize their inherent training dynamics, but also further improve overall efficiency, pruning performance, and representation quality simultaneously.
E.2 Taming-3DGS with AdamW-GS
| Pipeline | AdamW | All | Indoor | Outdoor | |||||||||||||
| -GS | PSNR | SSIM | LPIPS | Time | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||||
| Taming-3DGS | x | 27.386 | 0.796 | 0.258 | 0.668 | 0.620 | 7.44 | 31.025 | 0.918 | 0.205 | 0.357 | 0.329 | 24.479 | 0.698 | 0.299 | 0.916 | 0.853 |
| (multiplier) | ✓ | 27.537 | 0.799 | 0.254 | 0.646 | 0.575 | 5.27 | 31.268 | 0.920 | 0.200 | 0.357 | 0.313 | 24.552 | 0.703 | 0.297 | 0.877 | 0.785 |
| Taming-3DGS | x | 27.912 | 0.822 | 0.207 | 3.205 | 2.609 | 20.30 | 31.603 | 0.928 | 0.181 | 1.377 | 1.144 | 24.959 | 0.736 | 0.228 | 4.667 | 3.776 |
| (final count) | ✓ | 28.034 | 0.826 | 0.207 | 3.109 | 1.847 | 10.46 | 31.720 | 0.928 | 0.180 | 1.408 | 0.759 | 25.085 | 0.744 | 0.229 | 4.469 | 2.717 |
| Taming-3DGS-p | ✓ | 28.038 | 0.828 | 0.205 | 2.160 | 2.160 | 8.44 | 31.724 | 0.930 | 0.177 | 0.804 | 0.804 | 25.089 | 0.745 | 0.227 | 3.246 | 3.246 |
| (final count) | |||||||||||||||||
Taming-3DGS (Mallick et al., 2024) constrains the total number of primitives by restricting the number of new primitives added at each densification step, as defined by Eq. 12. During every densification step, Taming-3DGS computes a score for each primitive to determine its probability of being selected for densification. Based on these scores, a fixed number of primitives are sampled for densification. Additionally, Taming-3DGS incorporates further optimizations, such as enhanced parallelization. For detailed implementation and algorithmic design, please refer to (Mallick et al., 2024).
Where denotes the current number of densification steps, represents the total number of densification iterations, refers to the upper bound on the number of primitives after densification, and indicates the initial number of primitives.
Two operating modes are provided in Taming-3DGS444https://github.com/humansensinglab/taming-3dgs: final count and multiplier, which correspond to a higher and a lower upper bound on the final total primitive number , respectively. Experiments were conducted under both modes for Taming-3DGS as well as Taming-3DGS with AdamW-GS. DAR in this setting only includes opacity and scaling regularization as well, following the same configuration as in Sec. 4.4. Noise regularization is omitted because the growth rate in Taming-3DGS is constrained. Considering that AdamW-GS imposes stronger penalties on redundant primitives, leading to the generation of a large number of dead primitives that are rapidly pruned and thus reduce the total primitive count significantly, we propose to replace the conservative pruning strategy in Taming-3DGS with the pruning method used in vanilla 3DGS when integrating AdamW-GS. We denote this modified pipeline as Taming-3DGS-p. Related experimental results are summarized in Table 8.
As shown in Table 8, under the same constraint on the final , the variants trained with AdamW-GS consistently achieve better reconstruction quality compared with their original counterparts, while also providing additional advantages in training speed. Visual comparisons in Figure 6 further demonstrate that our method more faithfully preserves fine scene details. When a relatively large final primitive budget is given (the final count mode in Table 8), AdamW-GS reduces the total training time by nearly half (20.30 mins in the original Taming-3DGS vs. 10.46 mins for Taming-3DGS with AdamW-GS). Since AdamW-GS effectively penalizes redundant primitives at an early stage, we further incorporate the pruning strategy from vanilla 3DGS to construct Taming-3DGS-p. With our optimizer, this variant achieves additional GPU memory savings during training (3.205 million in the original Taming-3DGS vs. 2.160 million for Taming-3DGS with AdamW-GS).
E.3 Deformable Beta Splatting with AdamW-GS
| AdamW-GS | All | Outdoor | Indoor | ||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |
| x | 29.362 | 0.864 | 0.165 | 26.029 | 0.787 | 0.187 | 32.696 | 0.940 | 0.143 |
| ✓ | 29.643 | 0.871 | 0.158 | 26.108 | 0.798 | 0.175 | 33.178 | 0.945 | 0.140 |
Deformable Beta Splatting (DBS) (Liu et al., 2025a) shares a similar pipeline design and same loss with 3DGSMCMC (Kheradmand et al., 2024), as both are formulated under the Stochastic Gradient Langevin Dynamics framework. Unlike 3DGSMCMC, DBS replaces the Gaussian kernel with the Beta kernel and introduces the Spherical Beta function to better represent complex geometries and diverse appearance attributes. Consequently, the rendering process is reformulated as shown in Eq. 13.
Consistent with our experiments on 3DGSMCMC, we apply AdamW-GS to DBS, and the results are summarized in Table 9. Across eight of the nine scenes (excluding Treehill), we observe clear improvements in reconstruction quality when using AdamW-GS. For the Treehill scene, however, we find that DBS suffers from pronounced overfitting and triggers early stopping prematurely. Meanwhile, we observe that during the densification stage, has been closed early to zero, indicating that DAR cannot effectively exert its intended effect in this scenario. Our method may be less effective in such heavily overfitting scenario.
E.4 Vanilla 3DGS or 3DGSMCMC with AdamW-GS on OMMO Datasets
We provide results on the OMMO dataset (Lu et al., 2023), which contains large-scale outdoor scenes with long-range sequences. Our data processing follows the settings described in (Kheradmand et al., 2024) and its associated Github repository 555https://github.com/ubc-vision/3dgs-mcmc. However, the image preprocessing failed for scene #10; thus, we report the average results for the left scenes. All results are summarized in Table 10. Our method demonstrates effectiveness on long sequence datasets as well.
| AdamW-GS | 3DGSMCMC | 3DGS | ||||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |||||
| x | 30.359 | 0.925 | 0.135 | 1.960 | 1.673 | 30.040 | 0.914 | 0.154 | 1.878 | 1.640 |
| ✓ | 30.716 | 0.930 | 0.126 | 1.960 | 1.765 | 30.351 | 0.914 | 0.154 | 1.245 | 1.211 |
| Pipeline | All | Outdoor | Indoor | |||||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||||
| vanilla 3DGS | 27.506 | 0.815 | 0.216 | (3.098m) | 24.648 | 0.728 | 0.239 | (4.512m) | 31.080 | 0.925 | 0.189 | (1.331m) |
| + Only RSR (GS0) | 27.483 | 0.818 | 0.217 | -28.6% | 24.685 | 0.733 | 0.238 | -27.9% | 30.981 | 0.924 | 0.190 | -29.5% |
| +AdamW-GS (GS8) | 27.678 | 0.822 | 0.220 | -49.3% | 24.854 | 0.740 | 0.243 | -48.4% | 31.209 | 0.925 | 0.191 | -50.4% |
| +AdamW-GS (GS8) + ABE | 27.751 | 0.822 | 0.220 | -41.1% | 24.909 | 0.740 | 0.243 | -38.5% | 31.304 | 0.925 | 0.191 | -44.3% |
| +AdamW-GS (GS8) + Longer Densi | 27.715 | 0.824 | 0.218 | -48.4% | 24.857 | 0.744 | 0.240 | -44.6% | 31.288 | 0.925 | 0.190 | -53.2% |
Appendix F Exploration Strategies
In Sec. 4.1, we observe that Sparse Adam exhibits less explosive characteristic. Part of this comes from the stability described in Observation 1, and part is attributable to the inherent bad transportation property of 3DGS (Jung et al., 2024), which we briefly discussed in Sec. 4.4. This phenomenon causes a larger portion of primitives to remain in a stable state during optimization with Sparse Adam—neither triggering further primitive generation nor enabling continued improvements in scene reconstruction. As a result, we obtain a primitive group with a smaller but inferior reconstruction quality. In Sec. 5 (Extra Exploration is Necessary), we examine the necessity of additional exploration through experiments that introduce noise regularization. Here, exploration is deliberately to encourage primitives to undergo larger movements, which is consistent with the notion of bad transportation described in (Jung et al., 2024). We have given the discussion in Sec. 4.4.
Exploration in Current AdamW-GS
In Sec. 5, we argue that the current AdamW-GS improves exploration. This enhancement arises from several factors: (1) RSR. We provide additional ablation studies showing that, when using only Sparse Adam together with RSR, training the vanilla 3DGS pipeline on Mip-NeRF 360 yields reconstruction quality comparable to vanilla 3DGS with Adam. The corresponding results are reported in Table 11 under “+ Only RSR (GS0)”. (2) The influence of DAR around saddle points. The analysis in (Wang et al., 2025) indicates that many primitives become trapped near saddle points, preventing further effective optimization. The presence of DAR encourages these primitives to continue participating in optimization rather than stagnating. (3) Improved gradient flow after removing redundant primitives. When a large number of redundant primitives are eliminated, the gradient flow within primitive groups becomes more coherent and effective. (4) Noise regularization for outdoor scenes. For outdoor datasets, we intentionally introduce noise regularization to further encourage extra exploration.
Existing work provides only limited discussion of exploration strategies for the 3DGS pipeline. Noise regularization, while helpful in certain cases, has inherent limitations and is not applicable to all datasets. Nevertheless, we argue that developing extra exploration strategies is a promising direction, as such mechanisms encourage primitives to explore a broader region of the space. Building on our current work, we introduce two additional exploration strategies: Adaptive Bound-Expanding Split (Jung et al., 2024) and Densification Extending. (For outdoor scenes, the following experiments still employ noise regularization; we further demonstrate that these strategies can be used jointly.)
Adaptive Bound-Expanding Split (ABE-Split) divides each Gaussian into three, where the location of the third cross-region primitive is initialized using a constant factor proportional to the scene extent. The corresponding results are reported in Table 11 “+ AdamW-GS (GS8) + ABE”. ABE-Split consistently improves reconstruction quality across both indoor and outdoor scenes. In some cases, the gains are particularly clear—for example, on the Room scene from Mip-NeRF 360, the PSNR/SSIM increases from 31.500 dB / 0.920 (vanilla 3DGS) to 32.121 dB / 0.923.
Densification Extending: We extend the original densification phase from 15,000 iterations to 25,000 iterations, leveraging the densification mechanism. Thanks to the ability of our method to quickly penalize redundant primitives, this extension does not introduce the risk of memory explosion. The corresponding results are reported in Table 11 under “+ AdamW-GS (GS8) + Longer Densi”. Extending the densification phase further benefits indoor scenes and additionally reduces the number of active primitives.
Appendix G Failure Cases
Although AdamW-GS yields improved reconstruction quality—both quantitatively and qualitatively in terms of rendering visualization—compared with the original method, it only offers an optimization-level enhancement and does not fundamentally resolve the inherent limitations of the pipeline itself.
Severe artifacts in boundary regions with insufficient view coverage. As shown in Figure 7 a), boundary areas on walls exhibit substantial artifacts, floaters, and even incorrect colors due to the lack of sufficient multi-view constraints.
Insufficient refinement of background regions. These areas often remain under-optimized and may suffer from blurriness. We give an example in Figure 7 b).
Geometric inconsistency despite reasonable rendering results. Our current optimization does not incorporate any additional geometric priors. Obtaining the normalized depth map (Figure 7 c), Rendered Depth) using Eq. 14, it can be observed that many regions clearly violate geometric plausibility—for example, reflective books or shadowed chairs. Without explicit geometric priors, AdamW-GS cannot fully recover the correct underlying geometry. Nonetheless, certain improvements in geometry can be observed—for instance, the tree trunk region highlighted by the purple box in Figure 7 e).
Persisting floater issues when the original pipeline suffers from floaters. Even with AdamW-GS, floaters may remain, as illustrated in Figure 7 d).
Failure under severe overfitting. When the reconstruction pipeline significantly overfits the scene, AdamW-GS may break down. We encountered this issue when using DBS to reconstruct Treehill from the Mip-NeRF 360 dataset. Overfitting can prematurely trigger early stopping, and the sharp decreasing of leads to the failure of DAR.
where is the depth of the primitive in the corresponding camera coordinations.
Appendix H More about Related Work
Redundancy during Optimization
(1) Densification refinement: Several works optimize the densification rule (Fang & Wang, 2024; Mallick et al., 2024; Wang et al., 2025). SteepGS (Wang et al., 2025) mitigates redundancy through optimization-conditioned densification or alleviates sub-optimization by densification, but its reconstruction quality and pruning effectiveness remain inferior to learning-based methods. (2) Hand-crafted criterion: importance or redundancy criteria are designed to guide pruning after densification. LightGaussian (Fan et al., 2024) scores each Gaussian by opacity–transmittance–scale product times ray contributions. RadSplat (Niemeyer et al., 2025) uses the maximum alpha–transmittance product across views. (Papantonakis et al., 2024) propose the pipeline to prune primitives based on their degree of overlap. PUP-3DGS(Hanson et al., 2025) proposes a principled sensitivity pruning score which is computed as a second-order approximation of the reconstruction error on the training views. (3) Learning-based pruning that leverages learning-based strategies to adaptively remove redundant primitives, often via mask. Compact3DGS (Lee et al., 2024) introduces Gaussian masks with L1 regularization on active mask count. MaskGaussian (Liu et al., 2025b) models the existence of primitives, or measures the importance via learning. LP-3DGS(Zhang et al., 2024a) redesigns the masking function to leverage the Gumbel-Sigmoid method. (4) Optimization-related operations: Optimization together with additional opacity L1 regularization has become a common technique for redundancy removal, even extending beyond Efficient 3DGS tasks (Papantonakis et al., 2024; Lee et al., 2024; Kheradmand et al., 2024; Liu et al., 2025a; Svitov et al., 2024).
Optimizing redundancy and training efficiency constitute a subtopic of efficient 3DGS, corresponding to acceleration in inference and training (with partial overlap between the two). For inference, approaches include vector quantization (Lee et al., 2024), knowledge distillation (Fan et al., 2024), and entropy coding (Huang et al., 2025). For training, efforts target backward acceleration and loss redesign (Mallick et al., 2024). These topics fall beyond the scope of this work: for fair comparison, we maintain the same loss function and refrain from using additional compression techniques.
Optimizer and Regularization
The development of adaptive first-order optimizers can be traced back to RProp (Riedmiller & Braun, 1993). AdaGrad (Duchi et al., 2011) adapted the learning rate of features by estimated geometry and assigns larger learning rate to infrequent features. RMSProp (Hinton et al., 2012) further stabilized training by normalizing updates with an exponential moving average of squared gradients. Building on this, Adam (Kinga et al., 2015) incorporated momentum into RMSProp through an exponential average of gradients, quickly becoming the default optimizer for modern DNNs (Vaswani et al., 2017). Numerous refinements and extensions to Adam have since been proposed(You et al., 2019; Chen et al., 2023). Regularization has long been recognized as a key technique to improve model generalization and has also been employed as a tool to facilitate more effective optimization (Srivastava et al., 2014; Shalev-Shwartz & Ben-David, 2014; Andriushchenko et al., 2023; Brown et al., 2020; Radford et al., 2021). Importantly, Loshchilov & Hutter (2017) demonstrated that L2 regularization and weight decay are not equivalent, and further showed that weight decay provides a more appropriate formulation for Adam.
Optimization for 3DGS
Sparse Adam (Mallick et al., 2024) introduces asynchronous updates that improve efficiency, yet the insufficient recognition of update-step coupling results in degraded performance. 3DGS-LM (Höllein et al., 2025) proposes a tailored Levenberg-Marquardt optimizer for acceleration. However, it still relies on Adam for the initialization or densification. Although Adam has become a de facto indispensable optimizer for 3DGS, our understanding of its behavior in this context remains limited. A similar situation arises for regularization: despite the widespread use of various regularization techniques in 3DGS, there is currently little work examining the interplay between regularization and the optimizer, even though this has already been explored in the deep neural networks (Loshchilov & Hutter, 2017).
Appendix I Preparation for Decoupled Attribute Regularization
I.1 Hyperparameter Adjustment
This section discusses the influence of hyperparameters on regularization. For both opacity and scaling regularization in 3DGS-MCMC, a default value of 0.01 is adopted. Here, we vary only the opacity hyperparameter by scaling it up or down by a factor of 10, with the results summarized in Table 12. When increased to 0.1, the stronger regularization leads to severe degradation in reconstruction quality. Conversely, reducing it to 0.001 weakens the regularization effect to the extent that the primitive reallocation process is significantly disrupted, also resulting in quality loss. These results suggest that hyperparameters alone are insufficient to effectively regulate the strength of regularization
| Methods | Cite | Outdoor | Indoor | All | ||||||||||
| PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | ||||||
| MC2 | 0.01 | 25.160 | 0.754 | 0.212 | 3.92% | 31.546 | 0.930 | 0.183 | 4.74% | 27.998 | 0.832 | 0.199 | 4.28% | |
| MC101 | 0.1 | 17.379 | 0.400 | 0.604 | -77.8% | 29.957 | 0.906 | 0.220 | -40.7% | 22.969 | 0.625 | 0.433 | -59.3% | |
| MC102 | 0.001 | 24.585 | 0.731 | 0.220 | 5.18% | 31.342 | 0.930 | 0.179 | 8.1% | 27.588 | 0.819 | 0.202 | 6.28% | |
| AdamW o | MC103 | 0.1 | 24.298 | 0.714 | 0.242 | 5.79% | 30.687 | 0.921 | 0.193 | -2.08% | 27.138 | 0.806 | 0.220 | 2.83% |
| MC104 | 1 | 24.155 | 0.698 | 0.288 | 3.80% | 30.326 | 0.917 | 0.203 | -7.47% | 26.898 | 0.795 | 0.250 | 5.43% | |
| MC105 | 10 | 17.031 | 0.350 | 0.654 | -19.8% | 20.134 | 0.680 | 0.457 | -11.6% | 18.410 | 0.497 | 0.567 | -16.1% | |
| AdamW o | MC106 | 0.1 | 24.239 | 0.713 | 0.242 | 5.77% | 30.992 | 0.928 | 0.182 | 9.66% | 27.240 | 0.808 | 0.215 | 7.50% |
| + clip | MC107 | 1 | 24.381 | 0.713 | 0.260 | 5.08% | 30.850 | 0.922 | 0.194 | 8.76% | 27.256 | 0.806 | 0.231 | 6.71% |
| (max 10) | MC108 | 10 | 19.358 | 0.440 | 0.560 | 2.69% | 21.901 | 0.727 | 0.436 | 7.99% | 20.488 | 0.568 | 0.505 | 5.04% |
I.2 AdamW Style Decoupling
The AdamW-style decoupling implies an equal penalty applied to attributes, which can be formally expressed in Eq. 15 and Eq. 16. This formulation is equivalent to imposing a constant penalty on opacity (Rota Bulò et al., 2025). Within the 3DGS-MCMC framework, we apply AdamW-style decoupling to opacity and evaluate it under three hyperparameter settings representing different strengths. To disentangle the effect of magnitude, we also experiment with constraining the magnitude of the AdamW-style decoupling, as given in Eq. 17. The results are summarized in Table 12. Consistent with the findings in Appendix Sec. I.1, excessively large hyperparameters hinder convergence, while overly small values render the regularization ineffective. A key limitation of AdamW-style decoupling is that it enforces uniform penalties across all primitives, despite their varying importance.
I.3 Additional discussion regarding the formulation in Eq.8
A thorough decoupling—i.e., activation via the second moment of —is unreasonable: besides extra computational and memory cost, it reduces regularization to a trivial attribute-dependent penalty (e.g., higher opacity automatically incurs stronger regularization).
I.4 Discussion Related to Implicit Update
Ineffective when stable: Implicit updates allow primitives in invisible viewpoints to be updated based on their current moment. Our reasoning, grounded in experimental results, is as follows. When primitives are in a stable state—more precisely, at saddle points, which are abundant in 3DGS when no external perturbation is applied (Wang et al., 2025)—implicit updates are largely ineffective while incurring additional computational cost. In vanilla 3DGS, we implemented AIU (see GS4 and GS3 in Sec. K), and observed negligible improvement.
Less controllable when unstable: In contrast, when primitives are unstable, such as during the early training phase (e.g., densification, where newly generated primitives often have zero moment), or when the update step is relatively large, or when moment is rescaled, implicit updates can become problematic. We designed several experiments to highlight their uncontrollability. For instance, adopting AIU during densification in vanilla 3DGS with Sparse Adam can lead to more primitives but worse performance, as shown in GS6 in Appendix Sec. K, as even small update steps for invisible primitives introduce instability. Based on Obsidian 1, Sparse Adam tends to be stable, with which fewer primitives meet the gradient condition. AIU disrupts this stability: when primitives overfit certain views due to implicit updates, they induce larger gradients in other views, triggering unnecessary densification. The additional primitives do not necessarily improve rendering quality. In the Appendix Sec.K, we provide multiple AIU configurations: although in some cases the quality approaches that of vanilla 3DGS, the more common results are in worse quality, confirming the uncontrollable nature of implicit updates.
Another related phenomenon is the dead primitives or more reallocated primitives (see Sec.5.1). The comparison between GS1 and GS3 in Table 1 (or can be found in Appendix Sec. K) shows that more dead primitives are caused by Adam (0.232m vs. 0.048m). A natural conjecture is that when the update step, in a certain viewpoint, encourages primitives to become dead primitives, the implicit update accelerates this process, thereby producing more dead primitives. We provide some comparative experiments in Appendix Sec. K. (1) MC2 vs. MC3 in Appendix Sec. K, comparing the original 3DGSMCMC with Sparse Adam against the original 3DGSMCMC with Sparse Adam + AIU. (2) Similar comparisons can be found in GS3 and GS4, examining the effect of AIU on dead primitives when Sparse Adam is applied after densification ends. We observe that using AIU results in additional dead primitives. (3) Figure 3 likewise shows that adding AIU leads to more reallocated primitives during the densification stage of 3DGSMCMC. For certain scenes, such as Kitchen, the corresponding reconstruction quality decreases, further reflecting the increased number of dead primitives. This also indicates that the phenomenon is not part of a normal optimization process; in Sec. 5.1, we relate this behavior to over-effective regularization. (4) The comparison between MC6 and MC7 in Appendix Sec. K similarly shows that adding AIU after densification produces additional dead primitives as well; note that in our application RSR is only added during densification. (5) When training MaskGaussian with AdamW-GS, MaskGaussian will not suffer from the quality risk.
Appendix J Implementation Details
J.1 AdamW-GS
Specifically, each parameter of the optimizer can be expressed in the form:
Since both opacity and scaling are strictly positive, the gradient of the L1 regularization reduces to a constant value of .
For opacity:
For scaling:
For position:
For the other:
J.2 Hyperparameter Selection
For vanilla 3DGS(Kerbl et al., 2023) and 3DGS-MCMC(Kheradmand et al., 2024), all pipeline parameters follow the original settings, with noise parameters identical to those in 3DGS-MCMC. Operation for the opacity is similar as (Rota Bulò et al., 2025).
StSS in RSR
We fix the sampling interval at 100. Currently, StSS is manually set. StSS is used to sample the current set of primitives, followed by RSR rescaling. The configuration of StSS must take into account the generation and reallocation behavior in the pipeline, as these processes typically introduce additional primitives with zero state. When such zero-state primitives constitute a large proportion of the population, a smaller sampling ratio is generally preferred. In the main text, Figure 3 and Figure 2 respectively show the number of reallocated primitives for 3DGS-MCMC and the evolution of the primitive count under different scenes for 3DGS. Understanding these curves provides useful guidance for designing an effective StSS strategy.
In general, the differences between these schedules arise from two main factors: (1) variations in the underlying pipelines (e.g., 3DGS vs. 3DGS-MCMC), which naturally require distinct schedules; and (2) variations across scenes, where different schedules are needed to obtain the best performance on indoor versus outdoor datasets. The discrepancies introduced by the pipelines themselves are unavoidable. Nevertheless, for each pipeline we design a conservative schedule—denoted StSSGS1 and StSSMC1—under which both 3DGS and 3DGS-MCMC outperform their original configurations. For outdoor scenes, which typically involve higher primitive generation or reallocation rates, we find that a more aggressive schedule further improves performance. All StSS schedules used in this paper are illustrated in Figure 9. Appendix Sec.K summarizes the results obtained by applying different StSS configurations across various pipelines and scenes.
Scaling in RSR
The three existing or designed operations provide guidance for selecting the parameters and in RSR. (1) The states of newly generated primitives in 3DGS, or reallocated primitives in 3DGS-MCMC, can be viewed as naturally constrained by an RSR with . (2) To properly control the magnitude of the update step , it is necessary to impose the relation . (3) A very small is consistent with the design philosophy of DAR, as we rely on such a small value to rescale the second moment and thereby activate regularization. Therefore, when and are chosen to be sufficiently small while satisfying , the configuration is safe. We conducted a search over different combinations of and . For larger configurations—for example, —RSR becomes less effective, and some combinations even lead to gradient explosion. For smaller values of and , the performance differences are minor, and we find that generally performs best. In Appendix Sec.K, we report results for both and , cited as MC19 and MC18. Hence, we adopt as default in all experiments.
DAR
We use two DAR variants: opacity and scaling regularization. Regularization is constrained to the densification stage, as no clear benefit is observed in P-Op. This is consistent with the smaller max magnitude of adaptive gradients in P-Op (Figure 1). Opacity regularization starts after 3000 iterations, since a large percent of generated primitives initially have zero moment, motivating a delayed start.
For the two hyperparameters, we select values by considering activation derivatives, original settings, and Adam’s parameters. We set and, since scaling activations are two orders larger than opacity, . Comparative DAR experiments with and show that performs better (see experiments MC14 and MC17 in Appendix Sec. K). In our application, only the most significant digit of is retained and then scaled down by one order of magnitude.
For opacity, is set to 10, guided by the max magnitude of adaptive gradients in Figure 1. Without , over-effective regularization occurs (see MC9 in Appendix Sec. K). Smaller values (e.g., 5 or 1) reduce reallocated primitives, leading to slight drops in PSNR and SSIM, though the MCMC framework still explores effectively. For scaling, we adopt by analogy with opacity, which yields good results; scaling is less sensitive to (e.g., shows little change in reconstruction quality; see MC8–MC17 in Appendix Sec. K).
Appendix K All Results
More rendering visualization can be found in Fig.10.
K.1 Bicycle from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | reset | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | ✓ | - | 25.238 | 0.765 | 0.211 | 5.438 | 0.548 | - |
| GS2 | ✓ | - | - | - | - | ✓ | - | 25.142 | 0.752 | 0.235 | 4.052 | 0.100 | -25.5% |
| GS3 | Half | - | - | - | - | ✓ | - | 25.260 | 0.766 | 0.211 | 5.879 | 0.109 | 8.11% |
| GS0 | ✓ | - | - | - | - | ✓ | StSSGS1RSR | 25.337 | 0.770 | 0.213 | 3.929 | 0.038 | -28.4% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | ✓ | - | 25.284 | 0.766 | 0.210 | 5.832 | 0.176 | 7.24% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | ✓ | - | 25.256 | 0.766 | 0.210 | 5.651 | 0.285 | 3.91% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | ✓ | - | 25.145 | 0.756 | 0.227 | 5.245 | 0.105 | -3.54% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | ✓ | - | 25.166 | 0.755 | 0.228 | 5.026 | 0.103 | -7.57% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | ✓ | - | 25.166 | 0.757 | 0.226 | 5.290 | 0.102 | -2.72% |
| [0.1,1,7] | |||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | ✓ | - | 25.133 | 0.754 | 0.229 | 4.726 | 0.097 | -13.1% |
| [0.1,3] | |||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | ✓ | - | 25.116 | 0.756 | 0.226 | 5.343 | 0.103 | -1.74% |
| [0.1,3,7] | |||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | ✓ | - | 25.206 | 0.760 | 0.220 | 6.087 | 0.203 | 11.9% |
| GS9 | ✓ | - | - | - | ✓ | x | StSSGS2 | 25.481 | 0.776 | 0.219 | 3.116 | 0.013 | -42.7% |
| GS7 | ✓ | - | - | - | ✓ | x | StSSGS5 | 25.470 | 0.775 | 0.223 | 3.066 | 0.011 | -43.6% |
| GS10 | ✓ | - | - | - | ✓ | ✓ | StSSGS1 | 25.514 | 0.780 | 0.215 | 2.853 | 0.021 | -47.5% |
| GS8 | ✓ | - | - | - | ✓ | ✓ | StSSGS2 | 25.454 | 0.778 | 0.219 | 2.717 | 0.018 | -50.0% |
| cite | Sparse | AIU | ProAIU/ | RSR | / | / | PSNR | SSIM | LPIPS | /m | /m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 25.537 | 0.794 | 0.181 | 5.189 | 0.790 | -4.57% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 25.630 | 0.793 | 0.178 | 5.835 | 0.145 | 7.30% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 25.649 | 0.794 | 0.177 | 5.819 | 0.161 | 7.0% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 25.746 | 0.803 | 0.167 | 5.645 | 0.335 | 3.80% |
| MC5 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 24.930 | 0.798 | 0.163 | 5.764 | 0.216 | 5.99% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 25.298 | 0.801 | 0.161 | 5.789 | 0.191 | 6.45% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 25.779 | 0.804 | 0.170 | 5.734 | 0.246 | 5.44% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 25.777 | 0.803 | 0.166 | 5.722 | 0.258 | 5.22% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 25.539 | 0.795 | 0.180 | 5.189 | 0.791 | -4.57% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 25.766 | 0.803 | 0.167 | 5.715 | 0.265 | 5.09% |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 25.768 | 0.804 | 0.160 | 5.840 | 0.139 | 7.4% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 25.813 | 0.807 | 0.158 | 5.835 | 0.144 | 7.31% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 25.753 | 0.804 | 0.159 | 5.843 | 0.137 | 7.44% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 25.666 | 0.800 | 0.162 | 5.871 | 0.109 | 7.96% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 25.727 | 0.803 | 0.159 | 5.873 | 0.107 | 7.99% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 25.743 | 0.804 | 0.159 | 5.863 | 0.117 | 7.81% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 25.791 | 0.806 | 0.158 | 5.807 | 0.173 | 6.78% |
| MC17 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 25.781 | 0.806 | 0.158 | 5.807 | 0.173 | 6.78% |
| MC7 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 25.826 | 0.807 | 0.158 | 5.799 | 0.181 | 6.63% |
| MC8 | ✓ | - | - | StSSMC3 | DAR | 0.001/10 | DAR | /10 | 25.874 | 0.809 | 0.158 | 5.781 | 0.199 | 6.30% |
K.2 Flowers from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | reset | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | ✓ | - | 21.527 | 0.605 | 0.336 | 3.421 | 0.239 | - |
| GS2 | ✓ | - | - | - | - | ✓ | - | 21.463 | 0.598 | 0.346 | 2.737 | 0.030 | -19.9% |
| GS3 | Half | - | - | - | - | ✓ | - | 21.589 | 0.605 | 0.337 | 3.616 | 0.033 | 5.70% |
| GS0 | ✓ | - | - | - | - | ✓ | StSSGS1RSR | 21.698 | 0.610 | 0.336 | 2.506 | 0.011 | -26.7% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | ✓ | - | 21.598 | 0.606 | 0.336 | 3.581 | 0.056 | 4.67% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | ✓ | - | 21.585 | 0.605 | 0.336 | 3.518 | 0.098 | 2.83% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | ✓ | - | 21.488 | 0.600 | 0.343 | 3.198 | 0.032 | -6.51% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | ✓ | - | 21.535 | 0.600 | 0.342 | 3.153 | 0.032 | -7.83% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | ✓ | - | 21.563 | 0.601 | 0.342 | 3.240 | 0.032 | -5.29% |
| [0.1,1,7] | |||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | ✓ | - | 21.524 | 0.599 | 0.344 | 2.940 | 0.029 | -14.0% |
| [0.1,3] | |||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | ✓ | - | 21.548 | 0.601 | 0.342 | 3.330 | 0.003 | -2.66% |
| [0.1,3,7] | |||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | ✓ | - | 21.551 | 0.601 | 0.340 | 3.599 | 0.077 | 5.20% |
| GS9 | ✓ | - | - | - | ✓ | x | StSSGS2 | 21.762 | 0.609 | 0.343 | 2.026 | 0.004 | -40.7% |
| GS7 | ✓ | - | - | - | ✓ | x | StSSGS5 | 21.718 | 0.606 | 0.349 | 1.951 | 0.003 | -42.9% |
| GS10 | ✓ | - | - | - | ✓ | ✓ | StSSGS1 | 21.708 | 0.612 | 0.339 | 1.917 | 0.007 | -43.9% |
| GS8 | ✓ | - | - | - | ✓ | ✓ | StSSGS2 | 21.711 | 0.610 | 0.344 | 1.832 | 0.005 | -46.4% |
| cite | Sparse | AIU | ProAIU/ | RSR | / | / | PSNR | SSIM | LPIPS | /m | /m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 22.028 | 0.641 | 0.296 | 3.352 | 0.248 | -2.01% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 21.966 | 0.635 | 0.304 | 3.534 | 0.066 | 3.30% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 22.072 | 0.638 | 0.301 | 3.529 | 0.071 | 3.15% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 22.039 | 0.651 | 0.274 | 3.486 | 0.114 | 0.19% |
| MC5 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 21.750 | 0.650 | 0.271 | 3.520 | 0.08 | 2.89% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 21.511 | 0.646 | 0.275 | 3.524 | 0.076 | 3.01% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 22.195 | 0.649 | 0.292 | 3.510 | 0.090 | 2.60% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 22.038 | 0.650 | 0.277 | 3.524 | 0.076 | 3.01% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 22.017 | 0.641 | 0.295 | 3.352 | 0.248 | -2.01% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 21.288 | 0.588 | 0.349 | - | - | - |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 21.887 | 0.650 | 0.268 | 3.551 | 0.049 | 3.80% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 22.017 | 0.656 | 0.268 | 3.548 | 0.052 | 3.71% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 21.876 | 0.650 | 0.266 | 3.554 | 0.046 | 3.88% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 21.726 | 0.646 | 0.264 | 3.565 | 0.035 | 4.20% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 21.923 | 0.650 | 0.269 | 3.563 | 0.037 | 4.15% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 21.962 | 0.650 | 0.270 | 3.561 | 0.039 | 4.15% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 21.936 | 0.651 | 0.271 | 3.537 | 0.063 | 3.39% |
| MC17 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 22.034 | 0.653 | 0.270 | 3.539 | 0.061 | 3.44% |
| MC7 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 22.003 | 0.654 | 0.267 | 3.535 | 0.065 | 3.33% |
| MC8 | ✓ | - | - | StSSMC3 | DAR | 0.001/10 | DAR | /10 | 22.108 | 0.658 | 0.268 | 3.525 | 0.075 | 3.04% |
K.3 Garden from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | reset | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | ✓ | - | 27.334 | 0.865 | 0.108 | 5.614 | 0.318 | - |
| GS2 | ✓ | - | - | - | - | ✓ | - | 27.237 | 0.860 | 0.117 | 4.064 | 0.059 | -27.6% |
| GS3 | Half | - | - | - | - | ✓ | - | 27.409 | 0.866 | 0.107 | 5.878 | 0.059 | 4.70% |
| GS0 | ✓ | - | - | - | - | ✓ | StSSGS1RSR | 27.318 | 0.865 | 0.109 | 3.026 | 0.019 | -46.9% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | ✓ | - | 27.436 | 0.866 | 0.107 | 5.827 | 0.112 | 3.79% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | ✓ | - | 27.455 | 0.866 | 0.107 | 5.672 | 0.265 | 1.03% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | ✓ | - | 27.268 | 0.862 | 0.114 | 5.138 | 0.068 | -8.47% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | ✓ | - | 27.304 | 0.862 | 0.113 | 5.225 | 0.066 | -6.92% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | ✓ | - | 27.298 | 0.862 | 0.113 | 5.363 | 0.069 | -4.47% |
| [0.1,1,7] | |||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | ✓ | - | 27.234 | 0.861 | 0.114 | 4.652 | 0.061 | -17.1% |
| [0.1,3] | |||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | ✓ | - | 27.352 | 0.863 | 0.112 | 5.548 | 0.069 | -1.17% |
| [0.1,3,7] | |||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | ✓ | - | 27.263 | 0.862 | 0.112 | 5.877 | 0.162 | 4.68% |
| GS9 | ✓ | - | - | - | ✓ | x | StSSGS2 | 27.438 | 0.864 | 0.123 | 2.076 | 0.006 | -63.0% |
| GS7 | ✓ | - | - | - | ✓ | x | StSSGS5 | 27.590 | 0.864 | 0.127 | 1.983 | 0.005 | -64.7% |
| GS10 | ✓ | - | - | - | ✓ | ✓ | StSSGS1 | 27.370 | 0.863 | 0.122 | 2.145 | 0.016 | -61.8% |
| GS8 | ✓ | - | - | - | ✓ | ✓ | StSSGS2 | 27.170 | 0.862 | 0.124 | 2.075 | 0.012 | -63.0% |
| cite | Sparse | AIU | ProAIU/ | RSR | / | / | PSNR | SSIM | LPIPS | /m | /m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 27.862 | 0.879 | 0.094 | 5.466 | 0.433 | -2.63% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 27.889 | 0.878 | 0.096 | 5.785 | 0.115 | 3.04% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 27.896 | 0.878 | 0.095 | 5.770 | 0.13 | 2.77% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 28.032 | 0.883 | 0.089 | 5.627 | 0.273 | 0.23% |
| MC5 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 25.235 | 0.864 | 0.107 | 5.645 | 0.255 | 0.55% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 24.993 | 0.825 | 0.149 | 5.651 | 0.249 | 0.65% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 28.010 | 0.882 | 0.090 | 5.726 | 0.174 | 1.99% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 28.028 | 0.882 | 0.089 | 5.705 | 0.195 | 1.62% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 27.877 | 0.879 | 0.094 | 5.714 | 0.186 | 1.78% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 28.037 | 0.882 | 0.093 | 5.632 | 0.268 | 0.32% |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 28.087 | 0.884 | 0.088 | 5.752 | 0.148 | 2.24% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 28.065 | 0.884 | 0.088 | 5.740 | 0.160 | 2.24% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 28.072 | 0.884 | 0.088 | 5.756 | 0.144 | 2.52% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 28.021 | 0.883 | 0.089 | 5.807 | 0.093 | 3.43% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 28.045 | 0.883 | 0.088 | 5.805 | 0.095 | 3.40% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 28.018 | 0.883 | 0.088 | 5.796 | 0.104 | 3.24% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 28.081 | 0.884 | 0.088 | 5.725 | 0.175 | 1.97% |
| MC17 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 28.104 | 0.884 | 0.088 | 5.726 | 0.174 | 1.99% |
| MC7 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 28.109 | 0.884 | 0.088 | 5.714 | 0.186 | 1.78% |
| MC8 | ✓ | - | - | StSSMC3 | DAR | 0.001/10 | DAR | /10 | 28.144 | 0.885 | 0.088 | 5.687 | 0.213 | 1.30% |
K.4 Stump from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | reset | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | ✓ | - | 26.606 | 0.772 | 0.216 | 4.542 | 0.232 | - |
| GS2 | ✓ | - | - | - | - | ✓ | - | 26.511 | 0.762 | 0.233 | 3.751 | 0.026 | -17.4% |
| GS3 | Half | - | - | - | - | ✓ | - | 26.730 | 0.774 | 0.215 | 4.774 | 0.029 | 5.10% |
| GS0 | ✓ | - | - | - | - | ✓ | StSSGS1RSR | 26.624 | 0.782 | 0.210 | 3.644 | 0.016 | -19.7% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | ✓ | - | 26.710 | 0.773 | 0.215 | 4.928 | 0.094 | 8.49% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | ✓ | - | 26.759 | 0.774 | 0.214 | 4.596 | 0.270 | 1.18% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | ✓ | - | 26.534 | 0.766 | 0.226 | 4.974 | 0.029 | 9.51% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | ✓ | - | 26.475 | 0.764 | 0.227 | 5.010 | 0.030 | 10.3% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | ✓ | - | 26.611 | 0.768 | 0.224 | 4.811 | 0.026 | 5.94% |
| [0.1,1,7] | |||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | ✓ | - | 26.633 | 0.767 | 0.226 | 4.830 | 0.026 | 6.34% |
| [0.1,3] | |||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | ✓ | - | 26.630 | 0.767 | 0.224 | 5.165 | 0.026 | 13.7% |
| [0.1,3,7] | |||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | ✓ | - | 26.601 | 0.768 | 0.222 | 5.274 | 0.125 | 16.1% |
| GS9 | ✓ | - | - | - | ✓ | x | StSSGS2 | 27.175 | 0.796 | 0.211 | 2.894 | 0.014 | -36.3% |
| GS7 | ✓ | - | - | - | ✓ | x | StSSGS5 | 27.271 | 0.798 | 0.211 | 2.677 | 0.009 | -41.0% |
| GS10 | ✓ | - | - | - | ✓ | ✓ | StSSGS1 | 27.033 | 0.800 | 0.206 | 2.567 | 0.023 | -43.5% |
| GS8 | ✓ | - | - | - | ✓ | ✓ | StSSGS2 | 27.045 | 0.800 | 0.206 | 2.470 | 0.020 | -45.6% |
| cite | Sparse | AIU | ProAIU/ | RSR | / | / | PSNR | SSIM | LPIPS | /m | /m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 27.204 | 0.805 | 0.183 | 4.245 | 0.555 | -6.53% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 27.293 | 0.805 | 0.182 | 4.718 | 0.082 | 3.84% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 27.273 | 0.805 | 0.182 | 4.698 | 0.102 | 3.43% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 27.131 | 0.806 | 0.177 | 4.535 | 0.265 | -0.154% |
| MC5 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 20.854 | 0.687 | 0.284 | 4.621 | 0.179 | 1.73% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 25.509 | 0.782 | 0.195 | 4.615 | 0.185 | 1.60% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 27.064 | 0.805 | 0.179 | 4.620 | 0.180 | 1.71% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 27.172 | 0.807 | 0.176 | 4.631 | 0.169 | 1.93% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 27.170 | 0.796 | 0.190 | 4.673 | 0.127 | 2.88% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 26.943 | 0.796 | 0.191 | 4.620 | 0.180 | 1.71% |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 27.041 | 0.800 | 0.178 | 4.713 | 0.087 | 3.76% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 27.064 | 0.803 | 0.176 | 4.705 | 0.095 | 3.58% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 27.070 | 0.799 | 0.179 | 4.712 | 0.088 | 3.74% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 26.853 | 0.793 | 0.184 | 4.741 | 0.059 | 4.38% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 26.966 | 0.797 | 0.180 | 4.724 | 0.076 | 4.00% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 27.108 | 0.801 | 0.177 | 4.717 | 0.083 | 3.85% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 27.088 | 0.801 | 0.178 | 4.683 | 0.117 | 3.10% |
| MC17 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 27.101 | 0.802 | 0.178 | 4.681 | 0.119 | 3.06% |
| MC7 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 27.159 | 0.805 | 0.175 | 4.673 | 0.127 | 2.88% |
| MC8 | ✓ | - | - | StSSMC3 | DAR | 0.001/10 | DAR | /10 | 27.245 | 0.808 | 0.174 | 4.650 | 0.150 | 2.37% |
K.5 TreeHill from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | reset | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | ✓ | - | 22.534 | 0.633 | 0.326 | 3.547 | 0.228 | - |
| GS2 | ✓ | - | - | - | - | ✓ | - | 22.477 | 0.629 | 0.340 | 2.701 | 0.027 | -23.8% |
| GS3 | Half | - | - | - | - | ✓ | - | 22.558 | 0.634 | 0.327 | 3.820 | 0.031 | 7.69% |
| GS0 | ✓ | - | - | - | - | ✓ | StSSGS1RSR | 22.447 | 0.638 | 0.320 | 2.887 | 0.016 | -18.6% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | ✓ | - | 22.538 | 0.633 | 0.327 | 3.732 | 0.052 | 5.21% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | ✓ | - | 22.577 | 0.634 | 0.325 | 3.726 | 0.083 | 5.04% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | ✓ | - | 22.480 | 0.630 | 0.335 | 3.488 | 0.030 | -1.66% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | ✓ | - | 22.633 | 0.631 | 0.335 | 3.304 | 0.029 | -6.85% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | ✓ | - | 22.509 | 0.631 | 0.334 | 3.490 | 0.030 | -1.60% |
| [0.1,1,7] | |||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | ✓ | - | 22.537 | 0.630 | 0.335 | 3.131 | 0.026 | -11.7% |
| [0.1,3] | |||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | ✓ | - | 22.558 | 0.631 | 0.334 | 3.553 | 0.029 | 0.169% |
| [0.1,3,7] | |||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | ✓ | - | 22.484 | 0.631 | 0.331 | 4.074 | 0.080 | 14.8% |
| GS9 | ✓ | - | - | - | ✓ | x | StSSGS2 | 22.791 | 0.646 | 0.322 | 2.616 | 0.003 | -26.2% |
| GS7 | ✓ | - | - | - | ✓ | x | StSSGS5 | 22.691 | 0.644 | 0.328 | 2.553 | 0.002 | -28.0% |
| GS10 | ✓ | - | - | - | ✓ | ✓ | StSSGS1 | 22.905 | 0.650 | 0.322 | 2.335 | 0.006 | -34.1% |
| GS8 | ✓ | - | - | - | ✓ | ✓ | StSSGS2 | 22.891 | 0.651 | 0.325 | 2.231 | 0.005 | -37.1% |
| cite | Sparse | AIU | ProAIU/ | RSR | / | / | PSNR | SSIM | LPIPS | /m | /m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 22.894 | 0.655 | 0.310 | 3.329 | 0.37 | -6.14% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 23.023 | 0.659 | 0.300 | 3.621 | 0.079 | 2.08% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 23.012 | 0.658 | 0.302 | 3.612 | 0.088 | 1.83% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 22.858 | 0.663 | 0.284 | 3.555 | 0.145 | 0.22% |
| MC5 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 22.788 | 0.659 | 0.270 | 3.616 | 0.084 | 1.94% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 22.715 | 0.660 | 0.268 | 3.622 | 0.078 | 2.11% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 22.930 | 0.663 | 0.291 | 3.591 | 0.109 | 1.21% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 22.918 | 0.663 | 0.285 | 3.580 | 0.120 | 0.93% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 22.953 | 0.654 | 0.312 | 3.323 | 0.377 | -6.32% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | - | - | - | - | - | - |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 22.771 | 0.657 | 0.270 | 3.653 | 0.047 | 2.98% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 22.740 | 0.660 | 0.268 | 3.654 | 0.046 | 3.01% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 22.684 | 0.655 | 0.271 | 3.654 | 0.046 | 3.01% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 22.533 | 0.652 | 0.274 | 3.663 | 0.037 | 3.27% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 22.690 | 0.655 | 0.270 | 3.663 | 0.037 | 3.27% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 22.692 | 0.656 | 0.270 | 3.661 | 0.039 | 3.21% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 22.800 | 0.658 | 0.270 | 3.638 | 0.062 | 2.56% |
| MC17 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 22.806 | 0.658 | 0.270 | 3.639 | 0.061 | 2.59% |
| MC7 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 22.836 | 0.661 | 0.268 | 3.637 | 0.063 | 2.53% |
| MC8 | ✓ | - | - | StSSMC3 | DAR | 0.001/10 | DAR | /10 | 22.866 | 0.662 | 0.265 | 3.630 | 0.07 | 2.34% |
K.6 Room from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | - | 31.500 | 0.920 | 0.221 | 1.374 | 0.202 | - |
| GS2 | ✓ | - | - | - | - | - | 31.096 | 0.917 | 0.231 | 1.131 | 0.031 | -17.6% |
| GS3 | Half | - | - | - | - | - | 31.607 | 0.921 | 0.220 | 1.537 | 0.041 | 11.8% |
| GS0 | ✓ | - | - | - | - | StSSGS1RSR | 31.541 | 0.921 | 0.220 | 0.994 | 0.017 | -28.8% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | - | 31.719 | 0.921 | 0.220 | 1.496 | 0.111 | 8.87% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | - | 31.614 | 0.921 | 0.220 | 1.466 | 0.113 | 6.69% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | - | 31.389 | 0.919 | 0.228 | 1.447 | 0.034 | 5.31% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | - | 31.539 | 0.919 | 0.227 | 1.441 | 0.036 | 2.25% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | - | 31.379 | 0.918 | 0.227 | 1.473 | 0.037 | 7.20% |
| [0.1,1,7] | ||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | - | 31.476 | 0.919 | 0.228 | 1.342 | 0.034 | -2.32% |
| [0.1,3] | ||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | - | 31.509 | 0.919 | 0.226 | 1.504 | 0.036 | 9.46% |
| [0.1,3,7] | ||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | - | 31.259 | 0.917 | 0.228 | 1.549 | 0.066 | 12.7% |
| GS13 | ✓ | - | - | - | - | StSSGS3 | 31.768 | 0.920 | 0.227 | 0.617 | 0.006 | -55.1% |
| GS7/GS8 | ✓ | - | - | - | - | StSSGS1 | 31.741 | 0.921 | 0.222 | 0.624 | 0.005 | -54.6% |
| GS14 | ✓ | - | - | - | - | StSSGS4 | 31.753 | 0.920 | 0.222 | 0.632 | 0.005 | -54.0% |
| cite | Sparse | AIU | ProAIU/ | RSR | /maxo | /maxs | PSNR | SSIM | LPIPS | Na/m | Nd/m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 32.034 | 0.927 | 0.210 | 1.320 | 0.250 | -3.93% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 32.417 | 0.929 | 0.209 | 1.472 | 0.098 | 7.13% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 32.493 | 0.930 | 0.207 | 1.438 | 0.132 | 4.65% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 32.498 | 0.930 | 0.204 | 1.351 | 0.219 | -1.67% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 19.068 | 0.734 | 0.387 | 1.505 | 0.065 | 9.53% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 32.328 | 0.929 | 0.207 | 1.428 | 0.142 | 3.93% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 32.514 | 0.930 | 0.205 | 1.430 | 0.140 | 4.07% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 31.179 | 0.922 | 0.213 | 1.126 | 0.444 | -18.4% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 32.498 | 0.930 | 0.204 | 1.351 | 0.219 | -1.67% |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 32.600 | 0.933 | 0.196 | 1.536 | 0.034 | 11.7% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 32.542 | 0.933 | 0.197 | 1.535 | 0.035 | 11.7% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 32.611 | 0.933 | 0.197 | 1.536 | 0.034 | 11.7% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 32.317 | 0.932 | 0.197 | 1.546 | 0.024 | 12.5% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 32.547 | 0.933 | 0.195 | 1.528 | 0.042 | 11.2% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 32.586 | 0.933 | 0.196 | 1.528 | 0.042 | 11.2% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 32.647 | 0.933 | 0.197 | 1.515 | 0.055 | 10.2% |
| MC8/MC7 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 32.651 | 0.933 | 0.196 | 1.515 | 0.055 | 10.2% |
| MC21 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 32.647 | 0.933 | 0.197 | 1.512 | 0.058 | 10.0% |
K.7 Counter from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | - | 29.046 | 0.909 | 0.201 | 1.092 | 0.097 | - |
| GS2 | ✓ | - | - | - | - | - | 28.979 | 0.907 | 0.205 | 1.005 | 0.034 | -7.96% |
| GS3 | Half | - | - | - | - | - | 29.065 | 0.909 | 0.201 | 1.115 | 0.076 | 2.10% |
| GS0 | ✓ | - | - | - | - | StSSGS1RSR | 29.026 | 0.909 | 0.202 | 0.768 | 0.011 | -30.6% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | - | 29.060 | 0.909 | 0.201 | 1.124 | 0.067 | 2.93% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | - | 29.079 | 0.909 | 0.201 | 1.067 | 0.116 | -2.28% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | - | 28.947 | 0.908 | 0.204 | 1.163 | 0.035 | 6.50% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | - | 28.974 | 0.908 | 0.204 | 1.175 | 0.036 | 7.60% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | - | 29.010 | 0.908 | 0.203 | 1.199 | 0.036 | 9.97% |
| [0.1,1,7] | ||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | - | 28.972 | 0.908 | 0.203 | 1.118 | 0.033 | 2.38% |
| [0.1,3] | ||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | - | 28.998 | 0.908 | 0.202 | 1.221 | 0.036 | 11.81% |
| [0.1,3,7] | ||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | - | 28.974 | 0.908 | 0.202 | 1.257 | 0.058 | 15.1% |
| GS13 | ✓ | - | - | - | - | StSSGS3 | 29.084 | 0.907 | 0.204 | 0.579 | 0.005 | -46.9% |
| GS7/GS8 | ✓ | - | - | - | - | StSSGS1 | 29.112 | 0.907 | 0.206 | 0.535 | 0.004 | -51.0% |
| GS14 | ✓ | - | - | - | - | StSSGS4 | 29.077 | 0.907 | 0.203 | 0.535 | 0.003 | -51.0% |
| cite | Sparse | AIU | ProAIU/ | RSR | /maxo | /maxs | PSNR | SSIM | LPIPS | Na/m | Nd/m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 29.229 | 0.914 | 0.195 | 1.084 | 0.106 | -0.73% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 29.180 | 0.912 | 0.198 | 1.148 | 0.042 | 5.12% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 29.194 | 0.914 | 0.196 | 1.140 | 0.05 | 4.39% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 29.345 | 0.916 | 0.190 | 1.098 | 0.092 | 0.54% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 28.791 | 0.916 | 0.186 | 1.148 | 0.042 | 5.12% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 29.274 | 0.914 | 0.196 | 1.160 | 0.03 | 6.22% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 29.291 | 0.916 | 0.191 | 1.151 | 0.039 | 5.40% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 29.069 | 0.911 | 0.199 | 0.967 | 0.223 | -11.4% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 29.075 | 0.910 | 0.202 | 1.13 | 0.060 | 3.47% |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 29.476 | 0.919 | 0.184 | 1.172 | 0.018 | 7.32% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 29.495 | 0.919 | 0.183 | 1.171 | 0.019 | 7.23% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 29.470 | 0.919 | 0.184 | 1.143 | 0.047 | 4.67% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 29.450 | 0.919 | 0.183 | 1.181 | 0.009 | 8.15% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 29.455 | 0.919 | 0.182 | 1.172 | 0.018 | 7.32% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 29.486 | 0.919 | 0.183 | 1.171 | 0.019 | 7.23% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 29.475 | 0.919 | 0.184 | 1.163 | 0.027 | 6.50% |
| MC8/MC7 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 29.515 | 0.919 | 0.183 | 1.163 | 0.027 | 6.50% |
| MC21 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 29.506 | 0.919 | 0.184 | 1.161 | 0.029 | 6.31% |
K.8 Kitchen from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | - | 31.505 | 0.928 | 0.127 | 1.699 | 0.116 | - |
| GS2 | ✓ | - | - | - | - | - | 31.053 | 0.926 | 0.131 | 1.621 | 0.036 | -4.59% |
| GS3 | Half | - | - | - | - | - | 31.564 | 0.929 | 0.127 | 1.774 | 0.040 | 4.41% |
| GS0 | ✓ | - | - | - | - | StSSGS1RSR | 31.072 | 0.922 | 0.133 | 0.930 | 0.013 | -45.2% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | - | 31.531 | 0.929 | 0.127 | 1.721 | 0.089 | 1.29% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | - | 31.509 | 0.929 | 0.127 | 1.602 | 0.237 | -5.70% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | - | 31.110 | 0.926 | 0.130 | 1.766 | 0.035 | 3.94% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | - | 31.389 | 0.927 | 0.129 | 1.740 | 0.037 | 2.41% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | - | 30.997 | 0.926 | 0.129 | 1.769 | 0.035 | 4.12% |
| [0.1,1,7] | ||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | - | 31.330 | 0.927 | 0.129 | 1.724 | 0.035 | 1.47% |
| [0.1,3] | ||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | - | 31.355 | 0.927 | 0.129 | 1.772 | 0.036 | 4.29% |
| [0.1,3,7] | ||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | - | 31.404 | 0.928 | 0.128 | 1.782 | 0.069 | 4.88% |
| GS13 | ✓ | - | - | - | - | StSSGS3 | 31.770 | 0.927 | 0.131 | 0.692 | 0.005 | -59.3% |
| GS7/GS8 | ✓ | - | - | - | - | StSSGS1 | 31.728 | 0.927 | 0.132 | 0.667 | 0.004 | -60.7% |
| GS14 | ✓ | - | - | - | - | StSSGS4 | 31.741 | 0.926 | 0.133 | 0.667 | 0.004 | -60.7% |
| cite | Sparse | AIU | ProAIU/ | RSR | /maxo | /maxs | PSNR | SSIM | LPIPS | Na/m | Nd/m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 32.173 | 0.933 | 0.122 | 1.661 | 0.149 | -2.23% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 32.079 | 0.934 | 0.122 | 1.753 | 0.057 | 3.17% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 32.286 | 0.934 | 0.122 | 1.743 | 0.067 | 2.58% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 31.765 | 0.909 | 0.137 | 1.681 | 0.129 | -1.05% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 21.787 | 0.820 | 0.234 | 1.729 | 0.081 | 1.76% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 32.363 | 0.346 | 0.120 | 1.72 | 0.09 | 1.23% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 32.289 | 0.934 | 0.119 | 1.714 | 0.096 | 0.88% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 31.924 | 0.931 | 0.124 | 1.495 | 0.315 | -12.0% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 32.342 | 0.934 | 0.121 | 1.703 | 0.107 | 2.35% |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 32.570 | 0.937 | 0.117 | 1.765 | 0.045 | 3.88% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 32.397 | 0.935 | 0.118 | 1.763 | 0.047 | 3.76% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 32.443 | 0.936 | 0.118 | 1.767 | 0.043 | 4.00% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 32.458 | 0.935 | 0.118 | 1.786 | 0.024 | 5.12% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 32.466 | 0.936 | 0.117 | 1.766 | 0.044 | 3.94% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 32.173 | 0.935 | 0.118 | 1.766 | 0.044 | 3.94% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 32.521 | 0.936 | 0.117 | 1.756 | 1.756 | 3.35% |
| MC8/MC7 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 32.546 | 0.937 | 0.117 | 1.755 | 1.755 | 3.29% |
| MC21 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 32.298 | 0.936 | 0.117 | 1.753 | 0.057 | 3.17% |
K.9 Bosai from MipNerf-360
| cite | Sparse | AIU | ProAIU | noise | Ours | PSNR | SSIM | LPIPS | /m | /m | ||
| GS1 | x | - | - | - | - | - | 32.269 | 0.943 | 0.207 | 1.159 | 0.115 | - |
| GS2 | ✓ | - | - | - | - | - | 31.610 | 0.931 | 0.216 | 1.230 | 0.016 | 6.12% |
| GS3 | Half | - | - | - | - | - | 32.325 | 0.943 | 0.206 | 1.256 | 0.017 | 8.36% |
| GS0 | ✓ | - | - | - | - | StSSGS1RSR | 32.284 | 0.942 | 0.205 | 1.003 | 0.011 | -13.4% |
| GS4 | Half | [15,30] | 0.5 | [0.1][15] | - | - | 32.382 | 0.943 | 0.206 | 1.230 | 0.037 | 6.12% |
| GS15 | Half | [15,30] | 1.0 | [0.1][15] | - | - | 32.406 | 0.943 | 0.206 | 1.208 | 0.072 | 4.22% |
| GS16 | ✓ | [0.1,15] | 0.5 | [0.1][0.1] | - | - | 32.243 | 0.942 | 0.209 | 1.360 | 0.016 | 17.3% |
| GS17 | ✓ | [0.1,7] | 0.2 | [0.5][0.1] | - | - | 32.106 | 0.942 | 0.208 | 1.353 | 0.017 | 16.7% |
| GS18 | ✓ | [0.1,15] | 0.2 | [0.2,0.5,0.1] | - | - | 32.293 | 0.943 | 0.208 | 1.370 | 0.015 | 18.2% |
| [0.1,1,7] | ||||||||||||
| GS19 | ✓ | [0.1,15] | 0.2 | [0.5,0.1] | - | - | 32.260 | 0.941 | 0.209 | 1.283 | 0.015 | 10.6% |
| [0.1,3] | ||||||||||||
| GS5 | ✓ | [0.1,15] | 0.2 | [0.8,0.5,0.1] | - | - | 32.260 | 0.943 | 0.208 | 1.378 | 0.016 | 18.8% |
| [0.1,3,7] | ||||||||||||
| GS6 | ✓ | [0.1,30] | 1.0 | [0.1][0.1] | - | - | 32.251 | 0.942 | 0.208 | 1.428 | 0.034 | 23.2% |
| GS13 | ✓ | - | - | - | - | StSSGS3 | 32.409 | 0.942 | 0.204 | 0.807 | 0.004 | -30.3% |
| GS7/GS8 | ✓ | - | - | - | - | StSSGS1 | 32.251 | 0.942 | 0.205 | 0.747 | 0.004 | -35.5% |
| GS14 | ✓ | - | - | - | - | StSSGS4 | 32.279 | 0.941 | 0.206 | 0.735 | 0.003 | -36.5% |
| cite | Sparse | AIU | ProAIU/ | RSR | /maxo | /maxs | PSNR | SSIM | LPIPS | Na/m | Nd/m | |||
| MC1 | x | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 32.572 | 0.946 | 0.200 | 1.101 | 0.169 | -5.00% |
| MC2 | ✓ | - | - | - | L1 | 0.01/- | L1 | 0.01/- | 32.508 | 0.945 | 0.203 | 1.200 | 0.070 | 3.53% |
| MC3 | ✓ | [0.1,30] | 0.1/0.1 | - | L1 | 0.01/- | L1 | 0.01/- | 32.578 | 0.946 | 0.202 | 1.191 | 0.079 | 2.76% |
| MC4 | ✓ | [3,30] | 0.1/0.1 | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 32.742 | 0.947 | 0.198 | 1.130 | 0.140 | -2.50% |
| MC6 | ✓ | [26,30] | 0.01/1 | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 29.919 | 0.922 | 0.213 | 1.204 | 0.066 | 3.88% |
| MC18 | ✓ | - | - | StSSMC10 | L1 | 0.01/- | L1 | 0.01/- | 32.527 | 0.946 | 0.201 | 1.205 | 0.065 | 3.96% |
| MC19 | ✓ | - | - | StSSMC1 | L1 | 0.01/- | L1 | 0.01/- | 32.645 | 0.947 | 0.199 | 1.201 | 0.069 | 3.62% |
| MC20 | ✓ | - | - | StSSMC2 | L1 | 0.01/- | L1 | 0.01/- | 32.417 | 0.944 | 0.204 | 0.959 | 0.311 | -17.2% |
| MC9 | ✓ | - | - | StSSMC1 | DAR | 0.001/- | L1 | 0.01/- | 32.513 | 0.943 | 0.206 | 1.184 | 0.086 | 2.15% |
| MC10 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | L1 | 0.01/- | 33.006 | 0.950 | 0.191 | 1.221 | 0.049 | 5.34% |
| MC11 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | L1 | 0.01/- | 32.990 | 0.950 | 0.189 | 1.234 | 0.036 | 6.47% |
| MC12 | ✓ | - | - | StSSMC1 | DAR | 0.001/5 | L1 | 0.01/- | 32.990 | 0.950 | 0.189 | 1.237 | 0.033 | 6.72% |
| MC13 | ✓ | - | - | StSSMC1 | DAR | 0.001/1 | L1 | 0.01/- | 32.921 | 0.950 | 0.190 | 1.252 | 0.018 | 8.02% |
| MC14 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/10 | 32.916 | 0.950 | 0.189 | 1.233 | 0.037 | 6.38% |
| MC15 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | 0.001/1 | 32.906 | 0.950 | 0.189 | 1.232 | 0.038 | 6.29% |
| MC16 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /1 | 32.990 | 0.950 | 0.190 | 1.220 | 0.050 | 5.26% |
| MC8/MC7 | ✓ | - | - | StSSMC1 | DAR | 0.001/10 | DAR | /10 | 33.022 | 0.950 | 0.190 | 1.220 | 0.050 | 5.26% |
| MC21 | ✓ | - | - | StSSMC2 | DAR | 0.001/10 | DAR | /10 | 33.007 | 0.950 | 0.190 | 1.219 | 0.051 | 5.17% |